### v1

# Cross-Encoder Rediscovers a Semantic Variant of BM25

Abstract.
Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks. In this work, we investigate a Cross-Encoder variant of MiniLM to determine which relevance features it computes and where they are stored. We find that it employs a semantic variant of the traditional BM25 in an interpretable manner, featuring localized components: (1) Transformer attention heads that compute soft term frequency while controlling for term saturation and document length effects, and (2) a low-rank component of its embedding matrix that encodes inverse document frequency information for the vocabulary. This suggests that the Cross-Encoder uses the same fundamental mechanisms as BM25, but further leverages their capacity to capture semantics for improved retrieval performance. The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications.
explainability, neural retrieval models, IR axioms, mechanistic interpretability
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: ; July 13–18, 2025; Padua, Italyccs: Information systems Retrieval models and rankingccs: Computing methodologies Causal reasoning and diagnostics
1. Introduction
Information retrieval (IR) has long relied on foundational models like BM25, grounded in intuitive heuristics for relevance estimation. BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks. Its simplicity and interpretability have made it a cornerstone of traditional IR systems. Inspired by BM25’s success, earlier neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) were purposefully designed to emulate BM25’s principles. These models incorporated explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
However, the advent of transformer-based models has revolutionized the field of IR. These models, trained end-to-end on raw data, excel at extracting context-dependent semantic signals for ranking tasks. By leveraging the multi-attention mechanism and vast parameter spaces, transformers (Vaswani et al., 2023) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches. Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand how they assess relevance. This raises fundamental questions: how do these models assess relevance? Do they rely on established IR principles such as TF and IDF, or do they operate as vast, unstructured statistical approximators?
In this work, we build upon previous correlational studies by employing mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) methods to address these questions. These techniques have further advanced our understanding of the Transformer-based architectures (Vaswani et al., 2023) that underpin many modern NLP systems (Olsson et al., 2022; Geiger et al., 2021; Meng et al., 2024; Wang et al., 2022; Vig et al., 2020) and in IR, have been used to identify components responsible for encoding term frequency signals in bi-encoders (Chen et al., 2024).
In this paper, we deeply analyze a cross-encoder and validate prior research suggesting the presence of a BM25-like signal by causally demonstrating how a semantic variant of BM25 is precisely implemented within the components of a transformer architecture.
Specifically, our main findings are:
  • 4.14.3) We use path patching to identify attention heads that compute and process BM25-like components, including a semantic version of TF (soft-TF), term saturation, and document length. These heads send this information to another group of heads (Relevance Scoring Heads), which we hypothesize compute a BM25-style function by aggregating all these relevance signals.
  • 4.4): We find evidence that IDF information exists in a dominant low-rank vector of the embedding matrix and can be manipulated to control term importance.
  • 5): We confirm that Relevance Scoring Heads implement a BM25-style computation by defining a linear approximation of the cross-encoder’s relevance computation and evaluating its ability to reconstruct scores, confirming that our circuit understanding captures the core mechanism of relevance computation.
Table 1. Decomposition of BM25 into components and associated axioms. Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length which are tied to TF), we use alternative methods to isolate this component (§4.4).
BM25
Component
Axiom Perturbation Axiom-Based Diagnostic Dataset Examples
(b: baseline doc, p: perturbed doc)
TF (soft-TF) TFC1 (Fang et al., 2004): Prefer documents with more query term occurrences. Given a baseline document, we perturb it by appending one more selected query term. b: Quebec is a small city in Canada.
p: Quebec is a small city in Canada. Quebec.
STMC1 (Fang and Zhai, 2006): Prefer documents with semantically similar terms. Given a baseline document, we append a semantically similar term (the term with highest cross-encoder embedding similarity among 20 GPT-4o-generated candidates.) b: Quebec is a small city in Canada.
p: Quebec is a small city in Canada. Toronto.
IDF TDC (Fang et al., 2004): Prefer documents with more discriminative query terms. N/A N/A
k TFC2 (Fang et al., 2004): Additional occurrences yield smaller improvements. We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term. b: It is in Canada. It is a fun city.
p1: Quebec is in Canada. It is a fun city.
p2: Quebec is in Canada. Quebec is a fun city.
b LNC1 (Fang et al., 2004): Penalize longer documents for non-relevant terms. Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a random query in the base dataset, adding non-relevant information. b: Quebec is in Canada.
p1: Quebec is in Canada. I love papaya ice-cream at Abby’s.
p2: Quebec is in Canada. I love papaya ice-cream at Abby’s. Forever and Ever is a song by Julia.
By leveraging path patching for a sequential tracing of subcomponents on the cross-encoder’s term-matching path, we reverse-engineer a dominant portion of cross-encoder’s pathway to compute relevance. Such granular understanding reveals potential for model-editing, potentially enabling model performance enhance, controllability and personalization, and mitigate biases for NRMs.
2. Related Work
2.1. Interpreting Neural IR Models
In contrast to early neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) that were explicitly designed to be neural term matching models, recent pre-trained transformer-based ranking models achieve significant performance gains by capturing nuanced query-document interactions and forming rich contextual representations (Trabelsi et al., 2021; Guo et al., 2019). Although not explicitly designed to encode relevance concepts, these models have been found to learn relevance concepts such as TF (Chen et al., 2024), IDF (Choi et al., 2022), and even combinations of them such as BM25 (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). To better understand how these relevance signals are represented and utilized, researchers have turned to IR axioms (Câmara and Hauff, 2020; Rennings et al., 2019a; Völske et al., 2021) and constructed axiom-specific diagnostic datasets to test how well models rank document pairs in accordance with axiomatic constraints (Rennings et al., 2019b). Many studies also rely on correlational methods such as probing (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) to diagnose model behavior and assess their adherence to desired relevance properties. However, the underlying mechanisms of how these concepts are encoded remain unclear. This work aims to identify the specific components involved in this process and explain how they interact to determine relevance.
2.2. Mechanistic Interpretability
To better understand how information is encoded in models, mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) aims to reverse-engineer model behavior by analyzing the causal effects of components, providing stronger causal evidence than probing (Vig et al., 2020; Meng et al., 2024). Gaining traction in NLP, these methods have been used to uncover how LLMs perform tasks, such as indirect object identification (Meng et al., 2024) and greater-than computation (Hanna et al., 2024), and have recently been applied to IR to identify attention heads encoding term frequency signals (Chen et al., 2024).
Mechanistic interpretability methods rely on causal interventions, such as activation patching (or causal mediation analysis) where activations are swapped between input pairs to observe their effects on model behavior and localize specific mechanisms. Activation patching provides a solid foundation for initial analyses, allowing researchers to identify and isolate specific mechanisms within a model (Zhang and Nanda, 2023; Heimersheim and Nanda, 2024). For example, Chen et al. (2024) use activation patching with IR axioms to identify attention heads that encode a term frequency signal in a bi-encoder. However, activation patching has limitations as it only evaluates the indirect effect components have on the output. In contrast, path patching allows us to trace the computational path (circuit) from input to output, providing insights into how components sequentially interact (Wang et al., 2022; Goldowsky-Dill et al., 2023). In this work, we leverage path patching to reverse-engineer the computational path through which a NRM uses to compute relevance.
3. Methodology
Figure 1. Path patching methodology. Left: There are four forward passes: (1,2) Run model on baseline and perturbed inputs and cache activations. (3) To measure the effect of an upstream sender component (S) on a downstream receiver component (R), run the model on the baseline input, patch in S, freeze all other components, and cache the activation of R. (4) Run the model on the baseline input and patch in R. Right: Alternative visualization of step (3) using the residual stream. By allowing only the downstream receiver R to be recomputed when the sender S is patched, we effectively isolate the direct path from S to R, while preserving all other paths to R as they were in the baseline run.
In this section, we explain how we construct diagnostic datasets and use path patching to track how the model implements a semantic version of BM25.
3.1. Diagnostic Dataset Construction
Mapping BM25 components to IR axioms. IR axioms are derived from the same retrieval heuristics BM25 is composed of: TF, IDF, term saturation, and document length (Fang et al., 2004). Table 1 outlines the mapping from BM25 components to axioms and provides their formal definitions. Note that for TF, we extend this analysis of term matching to include semantic matches, where semantically similar terms contribute to the relevance score, based on the intuition about the semantic understanding capabilities of Transformer-based models (Trabelsi et al., 2021). We refer to this extension as soft-TF and utilize both TFC1 and STMC1 as the corresponding axioms.
Constructing minimal-pair-based diagnostic datasets. We construct a diagnostic dataset for each axiomatic component by perturbing documents to create input pairs. Our starting point is the base dataset from Chen et al. (Chen et al., 2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k query-document pairs. To create the diagnostic dataset for each axiom, we perturb each query-document pair in the base dataset in accordance with the formal axiomatic definition (Table 1). Each perturbed sample differs from the original only slightly, with the perturbation introducing an additional signal corresponding to the relevance concept we aim to analyze. Our perturbation strategies and examples of input pairs are shown in Table 1.
3.2. Path Patching
The intuition of path patching relies on the Transformer (Vaswani et al., 2023) architecture, where stacks of multi-headed attention and multi-layer perceptron (MLP) layers are connected by residual connections. These residual connections are additive functions, which essentially allow components to “read” and “write” information to a common residual stream.  (Team, 2021). The goal of path patching is to isolate which components “send” and “receive” information through the residual stream, and what information is being transmitted.
Path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) involves using minimal input pairs (we denote as baseline b and perturbed p), that differ only by a target behavioral signal to be evaluated (e.g., axiomatic property). Activations (e.g., residual stream output, attention head or layer output, MLP output) are iteratively copied from the perturbed input into the corresponding location in the baseline input to isolate the impact of that component on downstream computations. In this work, we focus on analyzing the roles of attention heads in relevance computation and leave analysis of MLPs for future investigation.
The path patching algorithm is an iterative backward process that starts with identifying which upstream components send important information to the logits (Figure 1). Specifically, it involves four forward passes through the model to identify which upstream (sender) components send information to the target (receiver) component (i.e., logits): (1) Run on the baseline input $x_{b}$ and cache activations. (2) Run on the perturbed input $x_{p}$ and cache activations. (3) Select the sender set $s$, the components whose activations are patched in, and the receiver set $r$, the components where the effect of patching $s$ is analyzed. Run the forward pass on $x_{b}$, patch in s, freeze all other activations, and recompute $r^{\prime}$, which is the same as $r$ from the $x_{b}$ run except for the direct influence from s to r. Cache $r^{\prime}$. (4) Run the model on $x_{b}$, and patch in $r^{\prime}$ values. Measure the difference in logits between $x_{b}$ and $x_{p}$ to quantify the effect of $s$ on $r$ in terms of passing the additional signal.
The effect of a patch is measured by the difference in logits (which in the case of cross-encoders, is equivalent to the difference in relevance scores). This algorithm is then iteratively repeated for each important upstream component. For more information, we refer the reader to Wang et al. (Wang et al., 2022) and Goldowsky-Dill et al. (Goldowsky-Dill et al., 2023).
We begin by applying path patching to track the path of term-matching, the most fundamental component of BM25. To carry this out, we use the TFC1 and STMC1 diagnostic datasets to identify the components responsible for encoding the relevance signal of an additional query term in a document, providing the foundation for our analysis of model behavior.
3.3. Model
We analyze ms-marco-MiniLM-L-12-v2 (Face, 2025) for our BERT-based cross-encoder, a model trained on the MS MARCO dataset, and chosen for its high evaluation performance the MS MARCO Passage Re-Ranking  (Nguyen et al., 2016) and TREC Deep Learning 2019 datasets (Craswell et al., 2020). The cross-encoder processes input x, consisting a query $q$ and a document $d$ in the format: <CLS> q1, q2, ... <SEP> d1, d2, ... <SEP>. After all transformer layers, the <CLS> token vector is passed to a classifier to produce the logits.
4. Semantic Scoring Circuit
Previous work in mechanistic interpretability has identified groups of attention heads whose behaviors generalize across tasks. For example, ”induction heads,” ”previous token heads,” and ”duplicate token heads” (Olsson et al., 2022) have been shown to play a role in various NLP tasks by completing sub-tasks within larger task-specific circuits. Identifying these heads is therefore crucial. Following the convention of naming head groups based on their mechanisms, we assign names to key components according to their specialized roles in document relevance computation. Figure 2 illustrates the uncovered relevance circuit, with details provided below:
  • Matching Heads generate a Matching Score by summing soft-TF values for each query term, while accounting for term saturation and document length effects.
  • Contextual Query Representation Heads distribute soft-TF information of the higher-IDF query tokens among all query tokens to strengthen the higher-IDF tokens’ representations.
  • IDF of vocabulary terms is stored in a dominant low-rank vector of the model’s Embedding Matrix.
  • Previously listed components send the four BM25 signals to Relevance Scoring Heads, which we hypothesize are summing Matching Score weighted by IDF in a BM25 manner.
Figure 2. We find that the cross-encoder rediscovers a semantic variant of BM25. Specifically, Matching Heads, Contextual Query Representation Heads, and the embedding matrix compute, process and send BM25-like components’ information to the Relevance Scoring Heads which finally compute the relevance score in a BM25-like manner.
As described in §3, we analyze the critical soft-TF component using path patching, tracing heads from ”back” to ”end” to align with its iterative nature. We first identify attention heads directly influencing logits and hypothesize they perform BM25-like computations by summing soft-TF weighted by IDF. Tracing further, we uncover Contextual Query Representation Heads and Matching Heads where soft-TF is originally computed. Using other diagnostic datasets, we find term saturation and document length signals intertwine with soft-TF at the Matching Heads. After tracking soft-TF, we apply SVD and intervention and localize IDF values in the first low-rank vector of the Embedding. We localize all BM25-like components in this section and verify that Relevance Scoring Heads perform BM25-like computations in the next section.
4.1. Relevance Scoring Heads
Figure 3. Path patching from individual attention heads to final logits reveals 10.1, 10.4, 10.7, 10.10 as the most critcal head in sending soft-TF information to [CLS] for ranking score computation for both TFC1 and STMC1 diagnostic datasets. Notably, these critical heads are shared between TFC1 and STMC1, exhibiting very similar importance values.
Information Flow. To identify components directly transmitting soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets.
Figure 3 highlights the most significant heads (i.e., those causing a $>30\%$ increase in ranking score) for the soft-TF-related axioms: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10. The high correlation in patching importance values (corr = 0.9996, p ¡ 0.001) and the fact that the four critical heads are shared between TFC1 and STMC1 suggests that the model treats exact and semantic term matches similarly.
Figure 4. Example of the attention pattern from [CLS] to query tokens in the Relevance Scoring Heads, illustrating how these heads process soft-TF for specific query tokens based on their IDF values.
Component Behavior. Upon closer inspection of the attention patterns of these heads, we observe that the [CLS] token selectively attends to query tokens and retrieves soft-TF based on each query token’s IDF value (Figure 4). Specifically, head 10.1 focuses predominantly on the highest-idf query token: the more uncommon a query token (i.e., the higher its IDF), the more the [CLS] token attends to it to capture its soft-TF information. Conversely, head 10.10 primarily attends to low-IDF tokens, while heads 10.7 and 10.4 focus on middle-to-high-IDF tokens.
To quantify this observation, we first compute the Pearson correlation between each head’s attention pattern and the query tokens’ IDF values. The results show strong correlations for heads 10.1 ($r=0.878$), 10.4 ($r=0.780$), and 10.7 ($r=0.798$), with $p<0.001$ for all, establishing a statistically significant relationship (threshold: $p<0.05$) between the attention values from [CLS] and the query tokens’ IDF values for 10.1, 10.4 and 10.7. However, Head 10.10 exhibits a weaker correlation ($r=0.206$). Then, we calculate the average IDF value weighted by each head’s attention distribution across query tokens to further validate this trend. Head 10.1 has the highest weighted average IDF (2.704), followed by heads 10.7 (1.983), 10.4 (2.171), and finally head 10.10 (1.457). This ranking quantitatively supports our observation and explains the weaker correlation between 10.10’s attention value with IDF values due to its negative association with them. Thus, each head covers different parts of the query based on IDF ranges, and together they send soft-TF information for all the query tokens to [CLS] together using a computation consisting of both soft-TF and IDF.
BM25-based Scoring Hypothesis. Considering that the mentioned computation resembles BM25, we hypothesize that the Relevance Scoring Heads perform a BM25-like summation of soft-TF values weighted by IDF of the query tokens. The model might attribute importance to the output of 10.1, 10.4 & 10.7, 10.10 from highest to lowest (e.g. controlled by different output projection matrix values) such that if a query token’s soft-TF is processed by 10.1, its soft-TF contributes the most to the relevance score. In this way, the query token’s soft-TF contribution to the relevance score is proportional to its IDF value. We verify this hypothesis in §5.
Figure 5. Example attention pattern of Matching Head 4.9: tokens attend most strongly to duplicates but also mildly to similar tokens.
Important Upstream Heads. Continuing our “backward pass,” we path patch to the relevance scoring heads to identify upstream components transmitting soft-TF information. This patching results for TFC1 and STMC1 show high correlation of 1.0 ( p ¡0.001) and reveals a shared set of attention heads, divided into two functional groups: (1) Contextual Query Representation Heads4.2) in middle layers, which reorganize soft-TF information among query tokens to construct a contextualized query representation. (2) Matching Heads4.3) in early layers, which compute Matching Scores (soft-TF weighted by term saturation and document length effects).
4.2. Contextual Query Representation Heads
The first group of heads that send information to Relevance Scoring Heads are termed contextual query representation heads: 8.10 and 9.11. These heads aggregate soft-TF information of the higher-IDF query tokens and distribute this information among all query tokens, strengthening their representations for the Relevance Scoring Heads’ final computation.
At the Relevance Scoring Heads, the [CLS] token retrieves soft-TF information from all query tokens. Thus, we analyze the attention patterns among query tokens in these two intermediary heads to understand how they modify the query token representation in the residual stream. The attention patterns reveal a distinct pattern: all query tokens in these heads consistently focus on one or two higher-IDF query tokens, with strong correlations to IDF values (9.11: $r=0.829$, 8.10: $r=0.781$ ) with both $p<0.001$. The evidence suggests that 8.10 and 9.11 act as Contextual Query Representation Heads: they distribute the soft-TF information of higher-IDF tokens across all query tokens. This ensures that when [CLS] token retrieves soft-TF information at Relevance Scoring Heads, each query token retains a stronger representation of the higher-IDF tokens’ soft-TF values.
Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads also receive soft-TF information from the Matching Heads.
4.3. Matching Heads
Matching Heads (0.8, 1.7, 2.1, 3.1, 4.9, 5.7, 5.9, 6.3, 6.5, 7.9, 8.0, 8.1, 8.8) compute and transmit a Matching Score, combining soft-TF, term saturation, and document length signals.
Component Behavior. Patching results reveal that Matching Heads transmit soft-TF information to the Relevance Scoring Heads and Contextual Query Representation Heads. A key question, however, is whether these heads merely propagate soft-TF signals from earlier components or compute the soft-TF themselves. Analyzing their attention patterns indicates that they actively attend to semantic matches and compute the soft-TF. Figure 5 illustrates this behavior for head 1.7: all query and documents tokens are strongly attending duplicates and mildly attending to semantically-similar terms.
Qualitative examination of the 13 heads with positive patching importance reveals attention patterns similar to those in Figure 5. We quantitatively check whether the query tokens are in fact matching semantically by measuring the correlation between the attention values (from any token to any document token) and the tokens’ semantic similarity, approximated using cross-encoder’s embeddings. The results show that Matching Heads exhibit a significantly higher average Pearson correlation (corr = 0.500, p¡0.001) compared to other heads (corr = 0.132), confirming that these heads function as Matching Heads: at these heads, query tokens actively attend to semantically similar terms, with the attention value serving as an indicator of semantic similarity. Following this intuition, the sum of attention values from a query token to all document tokens can approximate the soft-TF of a query term.
To validate the proposed function of the Matching Heads, we mean-ablate these Matching Heads and observe that the TF perturbed samples’ average value drops from 5.146 to -4.394, and the STMC perturbed samples’ average value drops from -1.998 to -4.611. The drastic drop occurs because ablation deactivates all semantic-matching functions for all query terms.
Figure 6. For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases. This aligns with TFC2, which states that additional term occurrences yield smaller improvements.
Figure 7. Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
Matching Score Contains Term Saturation and Document Length Signals. The average Soft-TF Correlation Score of Matching Heads (0.500) indicates that their attention values from query tokens to document tokens encode not only semantic similarity but also additional signals. To examine whether these attention values capture term saturation or document length effects, we analyze their behavior under using two diagnostic datasets: (1) TFC2 (increasing occurrences of a query term), and (2) LNC1 dataset (increasing number of irrelevant sentences to simulate the isolated effect of longer document length).
We define $x_{\text{selected}}$ as the selected query term, which is the duplicate term in TFC2 samples shown in Table 1. Similarly, $x_{\text{others}}$ refers to document tokens that are not duplicates of any query token. For each case, we calculate the normalized sum of attention values from all query tokens to $x_{\text{selected}}$ and $x_{\text{others}}$, representing the “total semantic match” of the matched token or the average across unmatched tokens, respectively.
As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 6). On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity. Furthermore, for most heads, the TFC2 trend is preserved in the attention values. For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 7). Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes. The varying degrees of influence on the heads’ attention values shown in Figure 6 and 7 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects. Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
Important Upstream Components. Due to time constraint, we further path patch to the three most important Matching Heads (7.9, 4.9, 2.1) to track how soft-TF flows into them. We find that the embedding exhibits the most direct influence, while previous attention and mlp layers exhibit very little influence. This completes our tracing of soft-TF path, confirming that the Matching Heads are the original “generator” of soft-TF as query token attends to soft matches to compute the Matching Score.
4.4. IDF in the Embedding Matrix
In addition to soft-TF, IDF is a critical signal for Relevance Scoring Heads. Here, we use SVD and low-rank intervention to track where IDF is stored or computed.
IDF exists in the first low-rank component of the embedding matrix. Previously, Choi et al.  (Choi et al., 2022) provide correlational evidence of IDF information existing in the embedding matrix. We decompose the embedding space into orthogonal components using Singular Value Decomposition (SVD). Recall that the embedding matrix $W_{E}$ is structured such that each row corresponds to a word’s representation in the vocabulary. The columns of $W_{E}$ represent different dimensions of this vector space, capturing various latent features of the words.
SVD allows us to decompose a complex matrix into a sum of rank-1 matrices, ordered by their contribution to the overall structure of the matrix. A rank-1 matrix is the outer product of a column and row vector, capturing a single direction in the data. This decomposition enables us to represent $W_{E}$ in orthogonal directions, allowing us to analyze the dominant directions in an easily interpretable manner due to their rank-1 structure. Mathematically, the SVD of $W_{E}$ is expressed as:
$W_{E}=USV^{T}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$
Here, $\sigma_{i}$ are the singular values, which quantify the importance or strength of each rank-1 component; $u_{i}$ and $v_{i}$ left and right singular vectors, respectively, forming orthonormal bases that capture patterns in the row (token embeddings) and column spaces (feature dimensions); and $r$ is the rank of $W_{E}$, representing the number of non-zero singular values and independent directions in the matrix. By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the Embedding.
Notably, we find that the rank-1 vector with the highest singular value, we call $U_{0}$, has a large negative pearson correlation ($-71.36\%$) with the IDF value from the MS-MARCO train set (Nguyen et al., 2016): higher values in this low-rank component for a given vocabulary token correspond to lower IDF values for that token. This suggests that IDF information is encoded prominently within the most significant low-rank component of the embedding matrix.
To validate this correlation, we perform an intervention in the next section to demonstrate that the IDF values from $U_{0}$ have a causal effect on downstream components (i.e., relevance scoring heads) and thus, the overall relevance computation.
Figure 8. $U_{0}$ editing example. Since $U_{0}$ is negatively correlated with IDF, increasing token 1’s (tok1) IDF, representing its importance in relevance computation, requires scaling its value in $U_{0}$ negatively. We transpose $U_{0}$ for visualization.
IDF intervention shows causal evidence that the model uses IDF in $U_{0}$ for relevance computation. Given previous understanding on $U_{0}$, we can interpret $U_{0}$ as a 1-D IDF dictionary, where $\text{idx}(q_{i})$ corresponds to the vocabulary index of $q_{i}$, mapping to its respective IDF value. Modifying the value at $\text{idx}(q_{i})$ in $U_{0}$ allows us to adjust the importance of the Matching Score for $q_{i}$ (Figure 8).
If the model uses the IDF values encoded in $U_{0}$, then modifying these values should result in corresponding changes to the ranking score. Specifically, according to the BM25-based Scoring Hypothesis in §4.1, there is a linear relationship between the IDF of a query token $q_{i}$ and the relevance score: increasing the IDF of $q_{i}$ should increase the relevance score. We have shown that $U_{0}$ is rank-1 and negatively correlated with IDF. If the model indeed uses the IDF values encoded in $U_{0}$, we can increase $\text{IDF}(q_{i})$ by decreasing qi’s value in U0, which would directly increase the relevance score. The converse should hold true for decreasing IDF.
One way to observe whether this is happening is to look at very controlled, minimal examples that we curate specifically to fully observe the effect. Using our base dataset, we create two documents, doc1 and doc2, containing distinct repeated tokens: first token (tok1) and second token (tok2), respectively (e.g., query: “computer science department number”, doc1: “computer computer computer”, doc2: “science science science”). Then, we can edit the IDF of tok1 and evaluate the IDF editing effect on the doc1 and doc2 relevance scores.
* We normalize the scores so that the average score of the unscaled doc1 is 0.
Figure 9. Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples. Larger increases in tok1’s IDF result in larger score increase for doc1. Bottom: Changes in scores for doc1 and doc2 when decreasing tok1’s IDF shows the inverse effect. Both increase and decrease has ranges where the monotonic trend reverses a little which suggests that there are optimal range for editing both ways.
Figure 9 offers support for the direct causal relationship: across most scaling factors, larger increase (or decrease) in tok1’s IDF result in larger score increase (or decrease, respectively) for doc1. The largely monotonic trend not only offers causal evidence for the model’s use of $U_{0}$ to acquire IDF values but also shows preliminary evidence supporting that the model computes a sum of soft-TF weighted by IDF values (§4.1).
Furthermore, the discovery of the low-rank nature of the IDF signals enables targeted interventions. Our intervention experiment demonstrates that modifying IDF values allows for precise control over the importance of specific terms within the cross-encoder.
5. Validation of BM25-like Computation
After we established where all the BM25-like components were initially computed or stored and how they were passed to Relevance Scoring Heads, we now assess whether the Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1. To do this, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function. Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
5.1. Formalizing the Hypothesized Function
Specifically, as proposed in the BM-based Scoring Hypothesis, the relevance scoring heads appear to compute a summation of Matching Scores weighted by IDF. The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function. If this hypothesis holds—that these components interact in a BM25-like manner—then the semantic scoring circuit can be expressed as a linear function:
$\sum_{i=1}^{N}\text{linear\_combo}\big{(}-U_{0}(q_{i}),\text{MS}_{\text{total}}(q_{i},d_{j}),[-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})]\big{)}$
where the components of the linear combination are defined as follows:
  1. (1)
    $U_{0}(q_{i})$: The value of $q_{i}$ in the $U_{0}$ vector, representing the model’s interpretation of $q_{i}$’s IDF.
  2. (2)
    $\text{MS}_{\text{total}}(q_{i},d_{j})$: The total Matching Score, computed as the sum of Matching Scores from individual heads ($\text{MS}_{H_{k}}$) weighted by learned weights $\alpha_{k}$. Each Matching Score represents the sum of attention values from $q_{i}$ to all document tokens:
    $\text{MS}_{\text{total}}(q_{i},d_{j})=\sum_{k=1}^{13}\alpha_{k}\cdot\text{MS}_{H_{k}}$
  3. (3)
    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, analogous to the product of IDF and TF in BM25.
Since the formalization contains not only the hypothesized computation function but also $U_{0}$ and $MS$ from earlier subparts of the circuit, the linear model is an effective representation of the circuit.
5.2. Assessing the Linear Model Fit
Given the linear representation, we can determine whether our hypothesis holds by comparing the linear model’s performance against the cross-encoder’s actual relevance scores.
We first train a linear regression model using our base dataset. As the number of coefficients in our linear regression scales quickly with each additional included query term, we apply a query length cutoff to five tokens in this initial experiment to first establish feasibility of representing the circuit in this manner. When performing forward passes on the samples, we extract the following for each query token: (1) the MS (Matching Score), calculated as the sum of the query token’s attention over document tokens from the 13 Matching Heads, and (2) the token’s corresponding value in $U_{0}$, obtained by applying SVD to the embedding matrix. These serve as the input features (“x”) for the linear regression model, while the cross-encoder’s ranking scores are used as the target values (“y”). We then split the data into an 80/20 train-test set to evaluate how well the linear representation of the Semantic Scoring Circuit predicts the cross-encoder’s relevance scores.
The linear regression model achieves a high pearson correaltion (corr = 0.8157, p ¡ 0.001) with ground-truth relevance scores, showing that it effectively captures the core components of the cross-encoder’s scoring mechanism in a simplified and interpretable form. This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p ¡ 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components more effectively capture the signals that the cross-encoder utilizes for ranking.
5.3. Generalization across IR datasets
We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept. Now, we ask the question, how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on? In this section, we extend our analysis to 12 different IR datasets [^1] Due to computational resources, we use a subset of BEIR datasets  (Thakur et al., 2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains. and investigate various query lengths.
Dataset. Since our linear model relies on fixed features based on query length, we stratify the dataset based on query length. To ensure sufficient samples, we aggregate samples from 12 datasets and only retain groups with more than 500 samples. We incorporate groups with query lengths from 1 to 22, which correponds to the 99.95th percentile of MSMARCO’s query length distribution to exclude outliers with excessively long queries that the model has rarely seen during training. This process results in a total of 19 groups, comprising 278194 samples, with an average group size of 14641.790 samples. While this approach may be unconventional for retrieval, our goal is to simply demonstrate consistency between the linear model’s relevance score and the cross-encoder to confirm the hypothesized function of the Relevance Scoring Heads.
First, as cross-encoders are typically used for re-ranking tasks, we follow the classic re-ranking setup by first retrieving a candidate set of documents (top 10) with BM25 over all queries across the test collection of the 12 datasets. These candidate sets are then split into train-test groups using an 80/20 ratio.
Experiments and Results. We train 22 linear regression models, one for each query length (1-22) to predict the cross-encoder’s relevance scores. We evaluate each model using: (1) Pearson Correlation: quantifies the correlation between predicted and actual relevance scores; (2) Spearman Rank Correlation: assesses the consistency of the ranked lists between predictions and cross-encoder outputs; (3) NDCG@10: measures ranking performance and ensures no significant performance discrepancies. We include BM25 and a randomized set of linear regression features as baselines.
Table 2 shows the results, and we report median values to account for observable skewness caused by outliers in the data. We observe a Pearson correlation of ranking scores with a median of 0.8401 (median p ¡ 0.001), a Spearman rank correlation of 0.7619 (median p = 0.072), and an 88.4% alignment with cross-encoder performance in terms of NDCG@10.
The experimental results confirm the hypothesized function of the Relevance Scoring Heads (§4.1). Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s performance shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
Table 2. Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
Model Pair Median Mean $\pm$ SD
Pearson Correlation
Cross-encoder & SemanticBM 0.8401 0.8301 $\pm$ 0.0314
Cross-encoder & Random Features 0.0005 0.0003 $\pm$ 0.0147
Cross-encoder & BM25 0.4570 0.4609 $\pm$ 0.0366
Spearman Rank Correlation
Cross-encoder & SemanticBM 0.7619 0.6629 $\pm$ 0.3137
Cross-encoder & Random Features 0.0000 0.0011 $\pm$ 0.3844
Cross-encoder & BM25 0.4643 0.3912 $\pm$ 0.3992
NDCG@10
Cross-encoder 0.5000 0.5097 $\pm$ 0.4110
SemanticBM 0.4307 0.4511 $\pm$ 0.3769
BM25 0.5000 0.5269 $\pm$ 0.4196
Random Features 0.3562 0.3498 $\pm$ 0.2986
6. Insights into Model Mechanisms
A two-stage process for query-document interactions. Our circuit analysis reveals that query-document interactions happen in two stages, bridging previous conflicting perspectives on how BERT cross-encoders form query representations. In the early to mid layers (0-8), matching heads attend to document terms that are exact or “soft” matches to query terms (§4.3). This first stage aligns with Qiao et al. (2019), who find that BERT focuses on exact query term matches or close paraphrases in the document, suggesting query-dependent document representations.
In mid to late layers (8-9), document representations are aggregated into query token representations by Contextual Query Representation heads (Section 4.2), consistent with Zhan et al. (2020) who argue that document tokens are largely query-independent. These two stages offer an explanation to previous conflicting perspectives on the role of query-document interactions: document representations initially have a query-specific relevance signal, but later, this signal is transferred to the query tokens. This dual-stage interaction process provides a more nuanced understanding of how query-document interactions evolve across cross-encoder layers.
Is (semantic) BM25 the whole story? While the Semantic Scoring circuit bears a strong resemblance to BM25, there also may be other factors influencing relevance computation that are out of the scope of investigation for this work. In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
Additionally, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores. There are a couple potential explanations for this: (1) the nonlinearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training (MacAvaney et al., 2022). We leave investigation of these avenues for future work.
Towards a holistic approach in axiomatic analysis. Previous research investigating NRMs’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones (Câmara and Hauff, 2020). Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves. BERT’s “rediscovery” of BM25 indicates that it leverages a combination of retrieval heuristics to compute relevance, whereas previous analysis treats axioms independently. Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.
7. Conclusion
In this work, we mechanistically uncovered the core components of relevance scoring pathway of an NRM, revealing their ability to leverage contextual representations to perform a semantic-variant of BM25. Our granular analysis paves way for future directions. For example, future work could build on our IDF editing technique (§4.4) to adjust term importance, offering more control over model behavior and leading to improvements in model robustness against adversarial attacks, personalization, and bias mitigation. In addition, our evidence about the centrality of IDF information in the model’s embedding matrix raises the question of how crucial non-embedding weights are for task generalization, and future work could investigate whether fine-tuning only the embedding matrix with domain-specific IDF values can achieve robust performance while being far more parameter efficient. Finally, in this work, we analyze the behavior of a single cross-encoder in order to deeply analyze its mechanisms, and future work should investigate their generalizability to other neural IR models. Overall, identifying universal components or architecture-specific mechanisms will advance interpretability and inform the design of controllable and efficient neural ranking systems for real-world applications.

### v2

@@
 Cross-Encoder Rediscovers a Semantic Variant of BM25
-Abstract.
+Abstract
 Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks.
@@
 The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications.
-explainability, neural retrieval models, IR axioms, mechanistic interpretability
-copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: ; July 13–18, 2025; Padua, Italyccs: Information systems Retrieval models and rankingccs: Computing methodologies Causal reasoning and diagnostics
-1. Introduction
-Information retrieval (IR) has long relied on foundational models like BM25, grounded in intuitive heuristics for relevance estimation.
+[^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25
+Cross-Encoder Rediscovers a Semantic Variant of BM25
+Meng Lu Brown University meng_lu@brown.edu                        Catherine Chen Brown University catherine_s_chen@brown.edu
+Carsten Eickhoff University of Tübingen carsten.eickhoff@uni-tuebingen.de
+1 Introduction
+Information retrieval (IR) is the subfield of NLP that aims to rank a collection of documents by their relevance to a specific query.
+Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (Robertson and Zaragoza, 2009), to estimate relevance.
 BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks.
-Its simplicity and interpretability have made it a cornerstone of traditional IR systems.
-Inspired by BM25’s success, earlier neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) were purposefully designed to emulate BM25’s principles.
-These models incorporated explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
-However, the advent of transformer-based models has revolutionized the field of IR.
-These models, trained end-to-end on raw data, excel at extracting context-dependent semantic signals for ranking tasks.
-By leveraging the multi-attention mechanism and vast parameter spaces, transformers (Vaswani et al., 2023) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches.
-Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand how they assess relevance.
+Its simplicity and inherent interpretability have made it a cornerstone of traditional IR systems.
+Inspired by BM25’s success, earlier neural IR models (Pang et al., 2016; Guo et al., 2016) were purposefully designed to emulate BM25’s principles.
+These models incorporate explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
+However, the advent of transformer-based models revolutionized the field of IR.
+These models, trained end-to-end on large numbers of query-document pairs (Nguyen et al., 2016; Thakur et al., 2021), excel at extracting context-dependent semantic signals for ranking tasks.
+By leveraging multi-headed attention and vast parameter spaces, transformers Vaswani et al.
+(2017) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches.
+Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand their internal mechanisms.
 This raises fundamental questions: how do these models assess relevance?
-Do they rely on established IR principles such as TF and IDF, or do they operate as vast, unstructured statistical approximators?
-In this work, we build upon previous correlational studies by employing mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) methods to address these questions.
-These techniques have further advanced our understanding of the Transformer-based architectures (Vaswani et al., 2023) that underpin many modern NLP systems (Olsson et al., 2022; Geiger et al., 2021; Meng et al., 2024; Wang et al., 2022; Vig et al., 2020) and in IR, have been used to identify components responsible for encoding term frequency signals in bi-encoders (Chen et al., 2024).
-In this paper, we deeply analyze a cross-encoder and validate prior research suggesting the presence of a BM25-like signal by causally demonstrating how a semantic variant of BM25 is precisely implemented within the components of a transformer architecture.
-Specifically, our main findings are:
-  • 4.14.3) We use path patching to identify attention heads that compute and process BM25-like components, including a semantic version of TF (soft-TF), term saturation, and document length.
-These heads send this information to another group of heads (Relevance Scoring Heads), which we hypothesize compute a BM25-style function by aggregating all these relevance signals.
-  • 4.4): We find evidence that IDF information exists in a dominant low-rank vector of the embedding matrix and can be manipulated to control term importance.
-  • 5): We confirm that Relevance Scoring Heads implement a BM25-style computation by defining a linear approximation of the cross-encoder’s relevance computation and evaluating its ability to reconstruct scores, confirming that our circuit understanding captures the core mechanism of relevance computation.
-Table 1.
-Decomposition of BM25 into components and associated axioms.
-Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length which are tied to TF), we use alternative methods to isolate this component (§4.4).
+Do they rely on established IR principles such as TF and IDF, or do they draw on entirely different expressions of relevance?
+In this work, we build upon previous correlational studies (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) by employing mechanistic interpretability methods (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) to address these questions.
+Concretely, we analyze a BERT-based IR model (i.e., a cross-encoder) to isolate multiple relevance signals beyond exact match TF.
+Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3).
+(2) We find evidence that IDF information largely already exists in the largest singular value of the embedding matrix and can be manipulated to later control term importance (§4.4).
+(3) We confirm that heads in later layers aggregate all these relevance signals in a BM25-style manner by defining a linear approximation of the hypothesized relevance computation and evaluating its ability to reconstruct cross-encoder’s predicted scores, confirming that our circuit captures the core mechanism of relevance computation (§5).
+Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer-based IR.
+2 Background and Related Work
+2.1 Axioms and BM25
 BM25
@@
 (b: baseline doc, p: perturbed doc)
-TF (soft-TF) TFC1 (Fang et al., 2004): Prefer documents with more query term occurrences.
+TF (soft-TF) TFC1 Fang et al.
+(2004): Prefer documents with more query term occurrences.
 Given a baseline document, we perturb it by appending one more selected query term.
@@
 p: Quebec is a small city in Canada. Quebec.
-STMC1 (Fang and Zhai, 2006): Prefer documents with semantically similar terms.
-Given a baseline document, we append a semantically similar term (the term with highest cross-encoder embedding similarity among 20 GPT-4o-generated candidates.) b: Quebec is a small city in Canada.
+STMC1 Fang and Zhai (2006): Prefer documents with semantically similar terms.
+Given a baseline document, we append a semantically similar term - the synonym with highest embedding cosine similarity with query term from 20 candidates generated by GPT-4o.
+b: Quebec is a small city in Canada.
 p: Quebec is a small city in Canada. Toronto.
-IDF TDC (Fang et al., 2004): Prefer documents with more discriminative query terms. N/A N/A
-k TFC2 (Fang et al., 2004): Additional occurrences yield smaller improvements.
+IDF TDC Fang et al. (2004): Prefer documents with more discriminative query terms. N/A N/A
+k TFC2 Fang et al.
+(2004): Additional occurrences yield smaller improvements.
 We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term.
 b: It is in Canada.
-It is a fun city.
-p1: Quebec is in Canada. It is a fun city.
-p2: Quebec is in Canada. Quebec is a fun city.
-b LNC1 (Fang et al., 2004): Penalize longer documents for non-relevant terms.
-Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a random query in the base dataset, adding non-relevant information.
+It is fun.
+p1: Quebec is in Canada. It is fun.
+p2: Quebec is in Canada. Quebec is fun.
+b LNC1 Fang et al.
+(2004): Penalize longer documents for non-relevant terms.
+Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a non-relevant query in the base dataset.
 b: Quebec is in Canada.
-p1: Quebec is in Canada. I love papaya ice-cream at Abby’s.
-p2: Quebec is in Canada. I love papaya ice-cream at Abby’s. Forever and Ever is a song by Julia.
-By leveraging path patching for a sequential tracing of subcomponents on the cross-encoder’s term-matching path, we reverse-engineer a dominant portion of cross-encoder’s pathway to compute relevance.
-Such granular understanding reveals potential for model-editing, potentially enabling model performance enhance, controllability and personalization, and mitigate biases for NRMs.
-2. Related Work
-2.1. Interpreting Neural IR Models
-In contrast to early neural IR models (e.g., MatchPyramid (Pang et al., 2016), DRMM (Guo et al., 2016)) that were explicitly designed to be neural term matching models, recent pre-trained transformer-based ranking models achieve significant performance gains by capturing nuanced query-document interactions and forming rich contextual representations (Trabelsi et al., 2021; Guo et al., 2019).
-Although not explicitly designed to encode relevance concepts, these models have been found to learn relevance concepts such as TF (Chen et al., 2024), IDF (Choi et al., 2022), and even combinations of them such as BM25 (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022).
-To better understand how these relevance signals are represented and utilized, researchers have turned to IR axioms (Câmara and Hauff, 2020; Rennings et al., 2019a; Völske et al., 2021) and constructed axiom-specific diagnostic datasets to test how well models rank document pairs in accordance with axiomatic constraints (Rennings et al., 2019b).
-Many studies also rely on correlational methods such as probing (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) to diagnose model behavior and assess their adherence to desired relevance properties.
-However, the underlying mechanisms of how these concepts are encoded remain unclear.
-This work aims to identify the specific components involved in this process and explain how they interact to determine relevance.
-2.2. Mechanistic Interpretability
-To better understand how information is encoded in models, mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) aims to reverse-engineer model behavior by analyzing the causal effects of components, providing stronger causal evidence than probing (Vig et al., 2020; Meng et al., 2024).
-Gaining traction in NLP, these methods have been used to uncover how LLMs perform tasks, such as indirect object identification (Meng et al., 2024) and greater-than computation (Hanna et al., 2024), and have recently been applied to IR to identify attention heads encoding term frequency signals (Chen et al., 2024).
-Mechanistic interpretability methods rely on causal interventions, such as activation patching (or causal mediation analysis) where activations are swapped between input pairs to observe their effects on model behavior and localize specific mechanisms.
-Activation patching provides a solid foundation for initial analyses, allowing researchers to identify and isolate specific mechanisms within a model (Zhang and Nanda, 2023; Heimersheim and Nanda, 2024).
-For example, Chen et al.
-(2024) use activation patching with IR axioms to identify attention heads that encode a term frequency signal in a bi-encoder.
-However, activation patching has limitations as it only evaluates the indirect effect components have on the output.
-In contrast, path patching allows us to trace the computational path (circuit) from input to output, providing insights into how components sequentially interact (Wang et al., 2022; Goldowsky-Dill et al., 2023).
-In this work, we leverage path patching to reverse-engineer the computational path through which a NRM uses to compute relevance.
-3. Methodology
-Figure 1.
-Path patching methodology.
+p1: Quebec is in Canada. Road not taken is written by Frost.
+p2: Quebec is in Canada. Road not taken is written by Frost. Happiness is a Butterfly.
+Table 1: Mapping of BM25 components to IR axioms.
+Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length, which are tied to TF), we use alternative methods to isolate it (§4.4).
+Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (Bruza and Huibers, 1994).
+For example, the TFC1 axiom (Fang et al., 2004) states that documents that contain more query terms should be ranked higher, and the TDC axiom (Fang et al., 2004) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores.
+These axioms provide a theoretical foundation for understanding and developing ranking functions by formalizing human-interpretable and intuitive notions of relevance into mathematical constraints.
+BM25 is a widely used probabilistic ranking function that exemplifies these axiomatic principles by ranking documents based on term frequency (TF), inverse document frequency (IDF), term saturation, and document length.
+It is defined as:
+$\sum_{t\in q}\text{IDF}(t)\cdot\frac{\text{TF}(t,d)\cdot(k_{1}+1)}{\text{TF}(t,d)+k_{1}\cdot(1-b+b\cdot\frac{|d|}{\text{avgdl}})}$ (1)
+where $q$ is the query, $d$ is the document, $t$ is a query term, $k_{1}$ and $b$ are hyperparameters controlling term saturation and length normalization, respectively.
+$|d|$ is the length of document $d$, and avgdl is the average document length in the corpus.
+Intuitively, BM25 computes a relevance score by multiplying each query term’s term frequency (TF) in the document with its inverse document frequency (IDF), capturing both how relevant and how informative the term is.
+It then sums these scores across all query terms with two adjustments: diminishing returns for repeated terms (term saturation via $k_{1}$), and downweighting for longer documents (length normalization via $|d|$), which prevents overly long documents with additional query term occurrences from being unfairly favored.
+As shown in Table 1, each component of BM25 aligns with a specific axiom.
+For instance, TFC1 reflects the TF factor, which favors documents containing more query terms.
+Framing BM25 in terms of axioms allows us to mechanistically analyze whether and how neural IR models internalize these principles when estimating relevance.
+2.2 Mechanistic Interpretability for IR
+Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022; Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022).
+To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Pearl, 2022), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification Meng et al.
+(2024) and greater-than computation Hanna et al.
+(2024).
+In the context of IR, Chen et al.
+(2024) apply activation patching (Vig et al., 2020; Meng et al., 2024) to identify attention heads responsible for exact term matching.
+We build on this work by decomposing BM25 into intuitive IR axioms and constructing diagnostic datasets to localize each component.
+Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023).
+This work is the first to causally uncover an internal circuit for the relevance estimation mechanism of neural IR architectures.
+3 Methodology
+3.1 Model
+While many early neural models were purposefully designed to emulate BM25’s principles, transformer-based models have revolutionized the field of IR.
+Among them, cross-encoders represent a specific architectural approach to neural ranking models that process query-document pairs jointly.
+Given input x in the format: <CLS> query <SEP> document <SEP>, the model is trained to optimize a binary classification task where the <CLS> token is passed to a classifier to determine whether the provided query is relevant to the document.
+In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (Face, 2025), a BERT-based cross-encoder, for its high performance on common IR benchmarks (Nguyen et al., 2016; Craswell et al., 2020).
+3.2 Diagnostic Datasets Construction
+To investigate if and how the model implements BM25, we map BM25 components to IR axioms (§2.1) and construct diagnostic datasets that isolate individual axiomatic components.
+We build off of Chen et al.
+(2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k web search query-document pairs.
+To create a diagnostic dataset for the remaining axioms, we perturb each query-document pair in this base dataset following their formal axiomatic definition (Table 1).
+Each perturbed sample differs minimally from the original, with the perturbation introducing an additional signal corresponding to the axiom.
+Our perturbation strategies and examples of input pairs are shown in Table 1.
+Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (Rau and Kamps, 2022).
+3.3 Path Patching on Diagnostic Datasets
+We first begin by tracking term matching, the core feature of classical IR models like BM25, by applying path patching on minimal pairs in TFC1 and STMC1 diagnostic datasets to uncover the soft-TF mechanism.
+For other components in BM25, we use these datasets to further inspect how they interact with the main soft-TF circuit.
+Specifically, we isolate soft-TF by selecting a query, a baseline document b, and a perturbed document p, where p introduces an additional semantic or lexical term frequency signal.
+We then substitute activations from p into b at specific model components, such as individual attention heads or MLP layers.
+If this substitution causes the model’s relevance score for b to shift toward that of p, we interpret it as causal evidence that the patched component encodes a term matching signal.
+The path patching algorithm proceeds iteratively in a backward fashion: it first identifies components directly responsible for changes in the output logits, then recursively traces upstream to uncover the full set of causally relevant components.
+This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix §A).
+4 Semantic Scoring Circuit
+Figure 1: Overview of relevance mechanisms in the model.
+In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (Figure 1):
+  • Matching Heads locate exact and semantic term matches in the document for each query term, while also encoding term saturation and document length signals.
+  • Query Contextualization Heads distribute the matching signal from higher-IDF query tokens to all query tokens.
+  • IDF of each term is stored in a dominant low-rank vector of the model’s embedding matrix.
+  • Relevance Scoring Heads aggregate query-term importance by combining the matching signal (soft-TF) and IDF for each query term in a manner similar to BM25.
+As described in §3, we begin by uncovering the core soft-TF mechanism via a “backward pass” of the model: path patching to the output logits, then progressively patching to earlier components to sequentially uncover important heads that compute and propagate soft-TF signals.
+After localizing the soft-TF mechanism, we apply Singular Value Decomposition (SVD) on the embeddings to investigate IDF.
+After isolating all BM25-like components in this section, we verify that Relevance Scoring Heads perform a BM25-like computation in §5.
+4.1 Relevance Scoring Heads
+Figure 2: Path patching identifies heads 10.1, 10.4, 10.7, and 10.10 as the most important carriers of soft-TF signals to [CLS] on both TFC1 and STMC1, with similar patching effects.
+Information Flow.
+To identify the components that directly transmit soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets.
+We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that the same heads produce similar effects on delivering soft-TF information into the final logits (see Appendix 9 for detailed patching results).
+Figure 2 displays the heads with the highest patching effects, which are consistent across both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10.
+Notably, these heads contribute a larger change in relevance scores for TFC1 compared to STMC1, suggesting that the model may prioritize exact matches over semantic ones.
+Figure 3: Example of the attention pattern from [CLS] to query tokens in the Relevance Scoring Heads, illustrating how these heads process soft-TF for specific query tokens based on their IDF values.
+Component Behavior.
+Qualitative analysis reveals that the [CLS] token selectively attends to query tokens, indicating the movement of soft-TF signals into its representation to prepare for the final relevance score in the model’s classification head.
+We find the heads distribute attention to different parts of the query.
+Specifically, 10.1 focuses on high-IDF query terms, 10.10 on low-IDF, and 10.7 and 10.4 on mid-to-high-IDF (Figure 3).
+We verify this distributive behavior by calculating the Pearson correlation between each head’s attention distribution and the IDF values of query tokens and find a moderately high average (r=0.67, p < 0.01).
+Comparison to BM25.
+Because these heads prepare the [CLS] token for the final scoring in the classification head, we call them Relevance Scoring Heads.
+They combine IDF and soft-TF signals per query term, resembling BM25’s term weighting mechanism.
+Thus, we hypothesize that the model combines the outputs of the Relevance Scoring Heads in a similar manner and verify this hypothesis in §5.
+Important Upstream Components.
+We path patch to the Relevance Scoring Heads’ value vectors to identify which upstream components transmit soft-TF signals.
+On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix C).
+4.2 Query Contextualization Heads
+Query Contextualization Heads (8.10 and 9.11) aggregate soft-TF signals of higher-IDF query tokens and distribute them across all query tokens, strengthening their representations for the Relevance Scoring Heads’ final computation.
+Component Behavior.
+At the Relevance Scoring Heads, the [CLS] token retrieves soft-TF from all query tokens.
+Thus, we analyze the attention patterns among query tokens in these two intermediary heads to understand how they modify the query token representation in the residual stream.
+We find that all query tokens in these heads consistently focus on one or two higher-IDF query tokens, with strong correlations to IDF values (9.11: $r=0.829$, 8.10: $r=0.781$ ) with both $p<0.001$.
+This suggests that 8.10 and 9.11 redistribute the soft-TF signals of higher-IDF tokens to contextualize the entire query for the Relevance Scoring Heads.
+Comparison to BM25.
+BM25 weights each query term independently, with no exchange of information between terms.
+In contrast, Query Contextualization Heads seem to learn to contextualize the entire query, redistributing soft-TF signals to amplify high-IDF tokens and dynamically reweight terms, before the final scoring.
+Further study of this learned contextualization is left for future work.
+Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (Appendix C).
+4.3 Matching Heads
+Figure 4: Example attention pattern of Matching Head 4.9: tokens attend most strongly to duplicates but also mildly to similar tokens.
+Alongside Query Contextualization Heads, a separate set of heads, named Matching Heads (0.8, 1.7, 2.1, 3.1, 4.9, 5.7, 5.9, 6.3, 6.5, 7.9, 8.0, 8.1, 8.8), in the model’s early and middle layers detect exact and semantic token matches (soft-TF) between the query and document and pass these signals to the Relevance Scoring Heads.
+Component Behavior.
+Qualitative inspection reveals that Matching Heads actively attend to both exact and semantically similar matches across the entire input sequence (i.e., query and document).
+As shown in Figure 4, head 1.7 exhibits this behavior clearly: both query and document tokens strongly attend to duplicated terms and, to a lesser extent, to semantically related terms.
+We compute the Pearson correlation between attention scores and semantic similarity for each query-document token pair, with semantic similarity measured as the cosine similarity between token embedding vectors.
+Matching Heads on average exhibit a significantly stronger correlation (r = 0.500, p < 0.001) than other heads (r = 0.132).
+This evidence suggests that Matching Heads employ an attention mechanism that scales proportionally with semantic proximity: higher semantic similarity corresponds to higher attention scores.
+Since each attention weight quantifies the strength of token matching, the sum of attention values from a query token to all document tokens effectively approximates the soft-TF of a query term.
+We validate the Matching Heads’ importance by mean-ablating them, and find that the average logits across both TFC1 and STMC1 perturbed samples decrease (TFC1: 5.146 to -4.394, STMC1: -1.998 to -4.611), indicating that these heads are critical for computing a semantic version of term matches.
+Additional Relevance Signals.
+Since BM25’s TF term takes two additional signals, term saturation and document length, into account, we also investigate whether they affect Matching Heads.
+To isolate these effects, we use two diagnostic datasets: (1) TFC2, which increases repeated query term occurrences to test for term saturation, and (2) LNC1, which adds irrelevant sentences to simulate the isolated effect of increased document length (Table 1).
+Because each dataset varies only in one factor, we track each Matching Head’s average attention across these controlled groups.
+Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (see Appendix D): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length.
+These findings suggest that Matching Heads integrate soft-TF with saturation and document-length signals rather than operating in isolation.
+We define this composite signal as the Matching Score, reflecting its incorporation of soft term frequency, term saturation, and document length effects, three core signals of BM25.
+Comparison to BM25.
+Matching Heads implement a semantic variant of the TF component in BM25 by identifying and weighting query-document token matches.
+Our TFC2 and LNC1 experiments show that they also go beyond simple match counting: their attention outputs are modulated in ways that mirror BM25’s additional relevance adjustments for term saturation and document length normalization.
+Thus, Matching Heads jointly capture all three core components of BM25, TF, term saturation, and length normalization, while also incorporating semantic similarity.
+Important Upstream Components.
+Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix C.2), confirming they generate soft-TF based on semantic similarity from the embeddings.
+This completes the backward tracing for the soft-TF circuit.
+4.4 IDF in the Embedding Matrix
+In addition to TF, IDF plays a critical role in BM25 by weighting each query token’s TF contribution, allowing the model to prioritize more uncommon terms.
+As shown in §4.1, IDF similarly emerges as a key signal for the Relevance Scoring Heads in our model.
+While Choi et al.
+(2022) provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
+SVD allows us to decompose a matrix into a sum of orthogonal rank-1 components, outer products of a column and row vector, ordered by their contribution to the overall matrix.
+Recall that the embedding matrix $W_{E}$ is structured such that each row corresponds to a word’s representation in the vocabulary, while the columns represent latent feature dimensions.
+By applying SVD to $W_{E}$, we can analyze the dominant directions to see what signals (e.g., IDF) drive the model’s behavior.
+Mathematically, the SVD of $W_{E}$ is expressed as:
+$W_{E}=USV^{T}=\sum_{i=1}^{r}\sigma_{i}\,u_{i}\,v_{i}^{T}$
+Here, $\sigma_{i}$ are the singular values, which quantify the importance or strength of each rank-1 component; $u_{i}$ and $v_{i}$ left and right singular vectors, respectively, forming orthonormal bases that capture patterns in the row (token embeddings) and column spaces (feature dimensions); and $r$ is the rank of $W_{E}$, representing the number of non-zero singular values and independent directions in the matrix.
+By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the embedding matrix.
+We find that the top singular vector $U_{0}$ is highly correlated (r = $-71.36\%$) with MS-MARCO IDF values Nguyen et al.
+(2016), the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
+Figure 5: $U_{0}$ editing example (transposed for visualization).
+Since $U_{0}$ is negatively correlated with IDF, increasing tok1’s IDF, representing its importance in relevance computation, requires negatively scaling its $U_{0}$ component.
+Causal Experiment.
+To validate this correlation, we perform an intervention to demonstrate that the IDF values from $U_{0}$ have a causal effect on downstream components (i.e., relevance scoring heads) and thus, the overall relevance computation.
+Given previous understanding on $U_{0}$, we can interpret $U_{0}$ as a 1-D IDF dictionary, where $\text{idx}(q_{i})$ corresponds to the vocabulary index of $q_{i}$, mapping to its respective IDF value.
+Modifying the value at $\text{idx}(q_{i})$ in $U_{0}$ allows us to adjust the importance of the Matching Score for $q_{i}$ (Figure 5).
+If the model uses the IDF values encoded in $U_{0}$, then modifying these values should result in corresponding changes to the ranking score.
+Specifically, according to the BM25-based Scoring Hypothesis in §4.1, there is a linear relationship between the IDF of a query token $q_{i}$ and the relevance score: increasing the IDF of $q_{i}$ should increase the relevance score.
+We have shown that $U_{0}$ is rank-1 and negatively correlated with IDF.
+If the model indeed uses the IDF values encoded in $U_{0}$, we can increase $\text{IDF}(q_{i})$ by decreasing $q_{i}$’s value in $U_{0}$, which would directly increase the relevance score.
+The converse should hold true for decreasing IDF.
+To test this, we design an experiment with controlled, minimal examples.
+Given a query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”).
+Then, we can edit the IDF of tok1 and measure the effect on the relevance scores of doc1 and doc2.
+* We normalize the scores so that the average score of the unscaled doc1 is 0.
+Figure 6: Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples.
+Increasing tok1’s IDF raises doc1’s score more than doc2’s.
+Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
+Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values (§4.1).
+Comparison to BM25.
+This completes the BM25 component set: IDF embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF.
+This is similar to BM25, where each query term’s TF is weighted by its IDF to prioritize informative terms.
+In the next section, we formally validate that Relevance Scoring Heads combine soft-TF and IDF in a BM25 manner to compute the final relevance score.
+5 Validation of BM25-like Computation
+To validate whether Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function.
+Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
+5.1 Formalizing the Hypothesized Function
+In §4.1, we hypothesize that Relevance Scoring Heads appear to compute a summation of Matching Scores weighted by IDF.
+The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function.
+If this hypothesis holds, that these components interact in a BM25-like manner, then the semantic scoring circuit can be expressed as a linear function:
+[…]
+$\quad-\,U_{0}(q_{i})\,\cdot\,\text{MS}_{\mathrm{total}}(q_{i},d_{j})\Bigr{)}$
+where the components of the linear combination are defined as follows:
+  1. 1.
+    $U_{0}(q_{i})$: The value of $q_{i}$ in the $U_{0}$ vector, representing the model’s interpretation of $q_{i}$’s IDF.
+  2. 2.
+    $\text{MS}_{\text{total}}(q_{i},d_{j})$: The total Matching Score, computed as the sum of Matching Scores from individual heads ($\text{MS}_{H_{k}}$) weighted by learned weights $\alpha_{k}$.
+Each Matching Score represents the sum of attention values from $q_{i}$ to all document tokens:
+    $\text{MS}_{\text{total}}(q_{i},d_{j})=\sum_{k=1}^{13}\alpha_{k}\cdot\text{MS}_{H_{k}}$
+  3. 3.
+    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
+By incorporating both the hypothesized computation function and the earlier components $U_{0}$ and $MS$, the linear model effectively represents the hypothesized circuit.
+5.2 Assessing the Linear Model Fit
+We test whether our hypothesis holds by comparing the linear model’s effectiveness against the cross-encoder’s actual relevance scores.
+First, we train a linear regression model using our base dataset, limiting queries to five tokens to keep the number of coefficients manageable.
+For each forward pass, we extract two features for each query token: (1) the MS (Matching Score), calculated as the sum of the query token’s attention over document tokens from the 13 Matching Heads, and (2) its value along the top singular vector in $U_{0}$ of the decomposed embedding matrix.
+These features form the input x to the linear regression model, while the cross-encoder’s relevance scores are the target y.
+Finally, we evaluate how well the linear representation of the Semantic Scoring Circuit predicts the cross-encoder’s relevance scores with an 80/20 train-test split.
+The linear regression model achieves a high Pearson correlation (r = 0.8157, p < 0.001) with ground-truth relevance scores, showing that it captures the core of the cross-encoder’s scoring mechanism in a simplified and interpretable form.
+This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking.
+The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix §E) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
+6 Discussion
+A two-stage process for relevance computation.
+Similar to Tenney et al.
+(2019), who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process.
+In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (Qiao et al., 2019).
+In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with Zhan et al.
+(2020), who argue that document tokens are largely query-independent.
+This dual-stage mechanism reconciles these two conflicting viewpoints by showing how query-specific signals migrate from document to query representations across layers and provides a more nuanced understanding of how relevance is computed.
+Controllable IR.
+In §4.4, we show how model editing can be used to upscale or downscale term importance by modifying the encoded IDF values in the embeddings, effectively introducing “tuning dials” for fine-grained control over model behavior.
+This opens the door to personalization, where models could be adapted to user- or domain-specific preferences without full retraining.
+Beyond personalization, these insights enable two additional downstream applications (preliminary experiments in Appendix §F).
+First, targeted model editing could mitigate unsafe or adversarial content without impairing the model’s ranking effectiveness.
+Second, aligning internal representations with ground-truth IDF scores could serve as a powerful initialization strategy for fine-tuning, leading to more efficient retrieval pipelines.
+7 Conclusion
+In this work, we mechanistically uncover the core components of the relevance scoring pathway in a BERT-based IR model, revealing how it leverages contextual representations to implement a semantic variant of BM25.
+Our fine-grained analysis lays the foundation for applying interventions to build more transparent and controllable models, enabling personalization, bias mitigation, and parameter-efficient adaptation.
+More broadly, identifying universal components or architecture-specific mechanisms contributes to the interpretability of IR models and informs the design of controllable, efficient ranking systems for real-world deployment.
+Limitations
+In this work, we focus on analyzing the behavior of a single cross-encoder in order to deeply analyze its mechanisms.
+Future work should investigate the generalizability of this behavior to other neural IR models with the same or different architectural base.
+Additionally, while the Semantic Scoring circuit bears a strong resemblance to BM25, there may also be other factors influencing relevance computation that are out of the scope of investigation for this work.
+In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022).
+However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
+Furthermore, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores.
+There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training MacAvaney et al.
+(2022).
+We leave the investigation of these avenues for future work.
+Finally, although we intend for our analysis to inform future work on efficiency, transparency, and performance improvements, like most research, it is possible that a malicious actor could apply these insights to produce IR models that are robust to adversarial use cases.
+Acknowledgment
+We thank the Health NLP Lab for valuable suggestions and discussions that improved this work.
+Appendix A Detailed Path Patching Methodology
+Figure 7: Path patching methodology.
 Left: There are four forward passes: (1,2) Run model on baseline and perturbed inputs and cache activations.
@@
 By allowing only the downstream receiver R to be recomputed when the sender S is patched, we effectively isolate the direct path from S to R, while preserving all other paths to R as they were in the baseline run.
-In this section, we explain how we construct diagnostic datasets and use path patching to track how the model implements a semantic version of BM25.
-3.1. Diagnostic Dataset Construction
-Mapping BM25 components to IR axioms.
-IR axioms are derived from the same retrieval heuristics BM25 is composed of: TF, IDF, term saturation, and document length (Fang et al., 2004).
-Table 1 outlines the mapping from BM25 components to axioms and provides their formal definitions.
-Note that for TF, we extend this analysis of term matching to include semantic matches, where semantically similar terms contribute to the relevance score, based on the intuition about the semantic understanding capabilities of Transformer-based models (Trabelsi et al., 2021).
-We refer to this extension as soft-TF and utilize both TFC1 and STMC1 as the corresponding axioms.
-Constructing minimal-pair-based diagnostic datasets.
-We construct a diagnostic dataset for each axiomatic component by perturbing documents to create input pairs.
-Our starting point is the base dataset from Chen et al.
-(Chen et al., 2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k query-document pairs.
-To create the diagnostic dataset for each axiom, we perturb each query-document pair in the base dataset in accordance with the formal axiomatic definition (Table 1).
-Each perturbed sample differs from the original only slightly, with the perturbation introducing an additional signal corresponding to the relevance concept we aim to analyze.
-Our perturbation strategies and examples of input pairs are shown in Table 1.
-3.2. Path Patching
-The intuition of path patching relies on the Transformer (Vaswani et al., 2023) architecture, where stacks of multi-headed attention and multi-layer perceptron (MLP) layers are connected by residual connections.
-These residual connections are additive functions, which essentially allow components to “read” and “write” information to a common residual stream.
-(Team, 2021).
-The goal of path patching is to isolate which components “send” and “receive” information through the residual stream, and what information is being transmitted.
-Path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) involves using minimal input pairs (we denote as baseline b and perturbed p), that differ only by a target behavioral signal to be evaluated (e.g., axiomatic property).
-Activations (e.g., residual stream output, attention head or layer output, MLP output) are iteratively copied from the perturbed input into the corresponding location in the baseline input to isolate the impact of that component on downstream computations.
-In this work, we focus on analyzing the roles of attention heads in relevance computation and leave analysis of MLPs for future investigation.
 The path patching algorithm is an iterative backward process that starts with identifying which upstream components send important information to the logits (Figure 7).
@@
 For more information, we refer the reader to Wang et al.
-(Wang et al., 2022) and Goldowsky-Dill et al.
-(Goldowsky-Dill et al., 2023).
+(2022) and Goldowsky-Dill et al.
+(2023).
 We begin by applying path patching to track the path of term-matching, the most fundamental component of BM25.
 To carry this out, we use the TFC1 and STMC1 diagnostic datasets to identify the components responsible for encoding the relevance signal of an additional query term in a document, providing the foundation for our analysis of model behavior.
-3.3. Model
-We analyze ms-marco-MiniLM-L-12-v2 (Face, 2025) for our BERT-based cross-encoder, a model trained on the MS MARCO dataset, and chosen for its high evaluation performance the MS MARCO Passage Re-Ranking  (Nguyen et al., 2016) and TREC Deep Learning 2019 datasets (Craswell et al., 2020).
-The cross-encoder processes input x, consisting a query $q$ and a document $d$ in the format: <CLS> q1, q2, ...
-<SEP> d1, d2, ...
-<SEP>.
-After all transformer layers, the <CLS> token vector is passed to a classifier to produce the logits.
-4. Semantic Scoring Circuit
-Previous work in mechanistic interpretability has identified groups of attention heads whose behaviors generalize across tasks.
-For example, ”induction heads,” ”previous token heads,” and ”duplicate token heads” (Olsson et al., 2022) have been shown to play a role in various NLP tasks by completing sub-tasks within larger task-specific circuits.
-Identifying these heads is therefore crucial.
-Following the convention of naming head groups based on their mechanisms, we assign names to key components according to their specialized roles in document relevance computation.
-Figure 2 illustrates the uncovered relevance circuit, with details provided below:
-  • Matching Heads generate a Matching Score by summing soft-TF values for each query term, while accounting for term saturation and document length effects.
-  • Contextual Query Representation Heads distribute soft-TF information of the higher-IDF query tokens among all query tokens to strengthen the higher-IDF tokens’ representations.
-  • IDF of vocabulary terms is stored in a dominant low-rank vector of the model’s Embedding Matrix.
-  • Previously listed components send the four BM25 signals to Relevance Scoring Heads, which we hypothesize are summing Matching Score weighted by IDF in a BM25 manner.
-Figure 2.
-We find that the cross-encoder rediscovers a semantic variant of BM25.
-Specifically, Matching Heads, Contextual Query Representation Heads, and the embedding matrix compute, process and send BM25-like components’ information to the Relevance Scoring Heads which finally compute the relevance score in a BM25-like manner.
-As described in §3, we analyze the critical soft-TF component using path patching, tracing heads from ”back” to ”end” to align with its iterative nature.
-We first identify attention heads directly influencing logits and hypothesize they perform BM25-like computations by summing soft-TF weighted by IDF.
-Tracing further, we uncover Contextual Query Representation Heads and Matching Heads where soft-TF is originally computed.
-Using other diagnostic datasets, we find term saturation and document length signals intertwine with soft-TF at the Matching Heads.
-After tracking soft-TF, we apply SVD and intervention and localize IDF values in the first low-rank vector of the Embedding.
-We localize all BM25-like components in this section and verify that Relevance Scoring Heads perform BM25-like computations in the next section.
-4.1. Relevance Scoring Heads
-Figure 3.
-Path patching from individual attention heads to final logits reveals 10.1, 10.4, 10.7, 10.10 as the most critcal head in sending soft-TF information to [CLS] for ranking score computation for both TFC1 and STMC1 diagnostic datasets.
-Notably, these critical heads are shared between TFC1 and STMC1, exhibiting very similar importance values.
-Information Flow. To identify components directly transmitting soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets.
-Figure 3 highlights the most significant heads (i.e., those causing a $>30\%$ increase in ranking score) for the soft-TF-related axioms: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10.
-The high correlation in patching importance values (corr = 0.9996, p ¡ 0.001) and the fact that the four critical heads are shared between TFC1 and STMC1 suggests that the model treats exact and semantic term matches similarly.
-Figure 4. Example of the attention pattern from [CLS] to query tokens in the Relevance Scoring Heads, illustrating how these heads process soft-TF for specific query tokens based on their IDF values.
-Component Behavior.
-Upon closer inspection of the attention patterns of these heads, we observe that the [CLS] token selectively attends to query tokens and retrieves soft-TF based on each query token’s IDF value (Figure 4).
-Specifically, head 10.1 focuses predominantly on the highest-idf query token: the more uncommon a query token (i.e., the higher its IDF), the more the [CLS] token attends to it to capture its soft-TF information.
-Conversely, head 10.10 primarily attends to low-IDF tokens, while heads 10.7 and 10.4 focus on middle-to-high-IDF tokens.
-To quantify this observation, we first compute the Pearson correlation between each head’s attention pattern and the query tokens’ IDF values.
-The results show strong correlations for heads 10.1 ($r=0.878$), 10.4 ($r=0.780$), and 10.7 ($r=0.798$), with $p<0.001$ for all, establishing a statistically significant relationship (threshold: $p<0.05$) between the attention values from [CLS] and the query tokens’ IDF values for 10.1, 10.4 and 10.7.
-However, Head 10.10 exhibits a weaker correlation ($r=0.206$).
-Then, we calculate the average IDF value weighted by each head’s attention distribution across query tokens to further validate this trend.
-Head 10.1 has the highest weighted average IDF (2.704), followed by heads 10.7 (1.983), 10.4 (2.171), and finally head 10.10 (1.457).
-This ranking quantitatively supports our observation and explains the weaker correlation between 10.10’s attention value with IDF values due to its negative association with them.
-Thus, each head covers different parts of the query based on IDF ranges, and together they send soft-TF information for all the query tokens to [CLS] together using a computation consisting of both soft-TF and IDF.
-BM25-based Scoring Hypothesis.
-Considering that the mentioned computation resembles BM25, we hypothesize that the Relevance Scoring Heads perform a BM25-like summation of soft-TF values weighted by IDF of the query tokens.
-The model might attribute importance to the output of 10.1, 10.4 & 10.7, 10.10 from highest to lowest (e.g.
-controlled by different output projection matrix values) such that if a query token’s soft-TF is processed by 10.1, its soft-TF contributes the most to the relevance score.
-In this way, the query token’s soft-TF contribution to the relevance score is proportional to its IDF value.
-We verify this hypothesis in §5.
-Figure 5. Example attention pattern of Matching Head 4.9: tokens attend most strongly to duplicates but also mildly to similar tokens.
-Important Upstream Heads.
-Continuing our “backward pass,” we path patch to the relevance scoring heads to identify upstream components transmitting soft-TF information.
-This patching results for TFC1 and STMC1 show high correlation of 1.0 ( p ¡0.001) and reveals a shared set of attention heads, divided into two functional groups: (1) Contextual Query Representation Heads4.2) in middle layers, which reorganize soft-TF information among query tokens to construct a contextualized query representation.
-(2) Matching Heads4.3) in early layers, which compute Matching Scores (soft-TF weighted by term saturation and document length effects).
-4.2. Contextual Query Representation Heads
-The first group of heads that send information to Relevance Scoring Heads are termed contextual query representation heads: 8.10 and 9.11.
-These heads aggregate soft-TF information of the higher-IDF query tokens and distribute this information among all query tokens, strengthening their representations for the Relevance Scoring Heads’ final computation.
-At the Relevance Scoring Heads, the [CLS] token retrieves soft-TF information from all query tokens.
-Thus, we analyze the attention patterns among query tokens in these two intermediary heads to understand how they modify the query token representation in the residual stream.
-The attention patterns reveal a distinct pattern: all query tokens in these heads consistently focus on one or two higher-IDF query tokens, with strong correlations to IDF values (9.11: $r=0.829$, 8.10: $r=0.781$ ) with both $p<0.001$.
-The evidence suggests that 8.10 and 9.11 act as Contextual Query Representation Heads: they distribute the soft-TF information of higher-IDF tokens across all query tokens.
-This ensures that when [CLS] token retrieves soft-TF information at Relevance Scoring Heads, each query token retains a stronger representation of the higher-IDF tokens’ soft-TF values.
-Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads also receive soft-TF information from the Matching Heads.
-4.3. Matching Heads
-Matching Heads (0.8, 1.7, 2.1, 3.1, 4.9, 5.7, 5.9, 6.3, 6.5, 7.9, 8.0, 8.1, 8.8) compute and transmit a Matching Score, combining soft-TF, term saturation, and document length signals.
-Component Behavior.
-Patching results reveal that Matching Heads transmit soft-TF information to the Relevance Scoring Heads and Contextual Query Representation Heads.
-A key question, however, is whether these heads merely propagate soft-TF signals from earlier components or compute the soft-TF themselves.
-Analyzing their attention patterns indicates that they actively attend to semantic matches and compute the soft-TF.
-Figure 5 illustrates this behavior for head 1.7: all query and documents tokens are strongly attending duplicates and mildly attending to semantically-similar terms.
-Qualitative examination of the 13 heads with positive patching importance reveals attention patterns similar to those in Figure 5.
-We quantitatively check whether the query tokens are in fact matching semantically by measuring the correlation between the attention values (from any token to any document token) and the tokens’ semantic similarity, approximated using cross-encoder’s embeddings.
-The results show that Matching Heads exhibit a significantly higher average Pearson correlation (corr = 0.500, p¡0.001) compared to other heads (corr = 0.132), confirming that these heads function as Matching Heads: at these heads, query tokens actively attend to semantically similar terms, with the attention value serving as an indicator of semantic similarity.
-Following this intuition, the sum of attention values from a query token to all document tokens can approximate the soft-TF of a query term.
-To validate the proposed function of the Matching Heads, we mean-ablate these Matching Heads and observe that the TF perturbed samples’ average value drops from 5.146 to -4.394, and the STMC perturbed samples’ average value drops from -1.998 to -4.611.
-The drastic drop occurs because ablation deactivates all semantic-matching functions for all query terms.
-Figure 6.
-For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases.
+Appendix B Activation Patching on Attention Heads
+Prior to path patching, we conduct an exploratory analysis to compare the similarity between exact and semantic matches.
+Figure 8 shows that the patching effect has a high correlation of 0.96, suggesting that the model employs highly similar components for both exact and soft matches.
+These initial results on the indirect effect of these components provide a foundation to investigate their direct effects further with path patching.
+Appendix C Detailed Path Patching Results
+C.1 Path Patching to Final Logits and Intermediate Heads
+We perform path patching on 2000 random query and document pairs from the STMC1 and TFC1 diagnostic datasets and collect the results.
+In this section, we present the results of path patching on TFC1 and STMC1.
+First, we patch to the final logits and identify the shared Relevance Scoring Heads (Figure 9).
+Next, we patch to these heads and discover the shared Query Contextualization Heads (Figures 1011).
+Finally, by patching to the Query Contextualization Heads, we identify the shared Matching Heads (Figure 12).
+Figure 8: Activation Patching Results for TFC and STMC diagnostic datasets shows a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
+Figure 9: Path Patching results from all attention heads to the final logits.
+From this, we identify that 10.1, 10.4, 10.7, 10.19, those causing > 30 % increase in ranking score, as most significant heads and focus our analysis on these four heads.
+Figure 10: Path Patching results from upstream attention heads to Relevance Scoring Heads.
+Figure 11: Path Patching results from upstream attention heads to Relevance Scoring Heads.
+Figure 12: Path Patching results from upstream attention heads to Query Contextualization heads.
+C.2 Path Patching to Matching Heads
+We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 13, 14, 15, 16).
+Path patching to the Matching Heads, aimed at tracking the flow of soft-TF information, reveals that at the beginning of the residual stream at layer 0, which corresponds to the embeddings, has the most direct influence on the Matching Heads.
+In contrast, there is minimal impact from both the attention and MLP layers.
+This supports the conclusion that the Matching Heads are the primary “generator” of soft-TF, attending to soft matches based on semantic similarity from the embeddings.
+Thus, we conclude our tracking of soft-TF information, with the Matching Heads identified as the most upstream heads.
+Figure 13: TFC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
+Figure 14: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
+Figure 15: STMC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
+Figure 16: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
+Appendix D Additional Details on Matching Heads and Additional Signals
+In this section, we provide additional details on the analysis presented in §4.3.
+Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence.
+The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, proving that these signals are discrete rather than binary.
+To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset.
+Figure 17 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals.
+This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
+Figure 17: For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases.
 This aligns with TFC2, which states that additional term occurrences yield smaller improvements.
-Figure 7.
-Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
+Figure 18: Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
 Matching Score Contains Term Saturation and Document Length Signals.
-The average Soft-TF Correlation Score of Matching Heads (0.500) indicates that their attention values from query tokens to document tokens encode not only semantic similarity but also additional signals.
+The average soft-TF correlation score for Matching Heads (0.500) indicates that their attention values from query tokens to document tokens capture not only semantic similarity but also additional signals.
 To examine whether these attention values capture term saturation or document length effects, we analyze their behavior under using two diagnostic datasets: (1) TFC2 (increasing occurrences of a query term), and (2) LNC1 dataset (increasing number of irrelevant sentences to simulate the isolated effect of longer document length).
@@
 On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity.
-Furthermore, for most heads, the TFC2 trend is preserved in the attention values.
-For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 7).
-Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes.
-The varying degrees of influence on the heads’ attention values shown in Figure 6 and 7 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects.
-Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
-Important Upstream Components.
-Due to time constraint, we further path patch to the three most important Matching Heads (7.9, 4.9, 2.1) to track how soft-TF flows into them.
-We find that the embedding exhibits the most direct influence, while previous attention and mlp layers exhibit very little influence.
-This completes our tracing of soft-TF path, confirming that the Matching Heads are the original “generator” of soft-TF as query token attends to soft matches to compute the Matching Score.
-4.4. IDF in the Embedding Matrix
-In addition to soft-TF, IDF is a critical signal for Relevance Scoring Heads. Here, we use SVD and low-rank intervention to track where IDF is stored or computed.
-IDF exists in the first low-rank component of the embedding matrix.
-Previously, Choi et al.
-(Choi et al., 2022) provide correlational evidence of IDF information existing in the embedding matrix.
-We decompose the embedding space into orthogonal components using Singular Value Decomposition (SVD).
-Recall that the embedding matrix $W_{E}$ is structured such that each row corresponds to a word’s representation in the vocabulary.
-The columns of $W_{E}$ represent different dimensions of this vector space, capturing various latent features of the words.
-SVD allows us to decompose a complex matrix into a sum of rank-1 matrices, ordered by their contribution to the overall structure of the matrix.
-A rank-1 matrix is the outer product of a column and row vector, capturing a single direction in the data.
-This decomposition enables us to represent $W_{E}$ in orthogonal directions, allowing us to analyze the dominant directions in an easily interpretable manner due to their rank-1 structure.
-Mathematically, the SVD of $W_{E}$ is expressed as:
-$W_{E}=USV^{T}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$
-Here, $\sigma_{i}$ are the singular values, which quantify the importance or strength of each rank-1 component; $u_{i}$ and $v_{i}$ left and right singular vectors, respectively, forming orthonormal bases that capture patterns in the row (token embeddings) and column spaces (feature dimensions); and $r$ is the rank of $W_{E}$, representing the number of non-zero singular values and independent directions in the matrix.
-By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the Embedding.
-Notably, we find that the rank-1 vector with the highest singular value, we call $U_{0}$, has a large negative pearson correlation ($-71.36\%$) with the IDF value from the MS-MARCO train set (Nguyen et al., 2016): higher values in this low-rank component for a given vocabulary token correspond to lower IDF values for that token.
-This suggests that IDF information is encoded prominently within the most significant low-rank component of the embedding matrix.
-To validate this correlation, we perform an intervention in the next section to demonstrate that the IDF values from $U_{0}$ have a causal effect on downstream components (i.e., relevance scoring heads) and thus, the overall relevance computation.
-Figure 8.
-$U_{0}$ editing example.
-Since $U_{0}$ is negatively correlated with IDF, increasing token 1’s (tok1) IDF, representing its importance in relevance computation, requires scaling its value in $U_{0}$ negatively.
-We transpose $U_{0}$ for visualization.
-IDF intervention shows causal evidence that the model uses IDF in $U_{0}$ for relevance computation.
-Given previous understanding on $U_{0}$, we can interpret $U_{0}$ as a 1-D IDF dictionary, where $\text{idx}(q_{i})$ corresponds to the vocabulary index of $q_{i}$, mapping to its respective IDF value.
-Modifying the value at $\text{idx}(q_{i})$ in $U_{0}$ allows us to adjust the importance of the Matching Score for $q_{i}$ (Figure 8).
-If the model uses the IDF values encoded in $U_{0}$, then modifying these values should result in corresponding changes to the ranking score.
-Specifically, according to the BM25-based Scoring Hypothesis in §4.1, there is a linear relationship between the IDF of a query token $q_{i}$ and the relevance score: increasing the IDF of $q_{i}$ should increase the relevance score.
-We have shown that $U_{0}$ is rank-1 and negatively correlated with IDF.
-If the model indeed uses the IDF values encoded in $U_{0}$, we can increase $\text{IDF}(q_{i})$ by decreasing qi’s value in U0, which would directly increase the relevance score.
-The converse should hold true for decreasing IDF.
-One way to observe whether this is happening is to look at very controlled, minimal examples that we curate specifically to fully observe the effect.
-Using our base dataset, we create two documents, doc1 and doc2, containing distinct repeated tokens: first token (tok1) and second token (tok2), respectively (e.g., query: “computer science department number”, doc1: “computer computer computer”, doc2: “science science science”).
-Then, we can edit the IDF of tok1 and evaluate the IDF editing effect on the doc1 and doc2 relevance scores.
-* We normalize the scores so that the average score of the unscaled doc1 is 0.
-Figure 9.
-Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples.
-Larger increases in tok1’s IDF result in larger score increase for doc1.
-Bottom: Changes in scores for doc1 and doc2 when decreasing tok1’s IDF shows the inverse effect.
-Both increase and decrease has ranges where the monotonic trend reverses a little which suggests that there are optimal range for editing both ways.
-Figure 9 offers support for the direct causal relationship: across most scaling factors, larger increase (or decrease) in tok1’s IDF result in larger score increase (or decrease, respectively) for doc1.
-The largely monotonic trend not only offers causal evidence for the model’s use of $U_{0}$ to acquire IDF values but also shows preliminary evidence supporting that the model computes a sum of soft-TF weighted by IDF values (§4.1).
-Furthermore, the discovery of the low-rank nature of the IDF signals enables targeted interventions.
-Our intervention experiment demonstrates that modifying IDF values allows for precise control over the importance of specific terms within the cross-encoder.
-5. Validation of BM25-like Computation
-After we established where all the BM25-like components were initially computed or stored and how they were passed to Relevance Scoring Heads, we now assess whether the Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1.
-To do this, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function.
-Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
-5.1. Formalizing the Hypothesized Function
-Specifically, as proposed in the BM-based Scoring Hypothesis, the relevance scoring heads appear to compute a summation of Matching Scores weighted by IDF.
-The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function.
-If this hypothesis holds—that these components interact in a BM25-like manner—then the semantic scoring circuit can be expressed as a linear function:
-$\sum_{i=1}^{N}\text{linear\_combo}\big{(}-U_{0}(q_{i}),\text{MS}_{\text{total}}(q_{i},d_{j}),[-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})]\big{)}$
-where the components of the linear combination are defined as follows:
-  1. (1)
-    $U_{0}(q_{i})$: The value of $q_{i}$ in the $U_{0}$ vector, representing the model’s interpretation of $q_{i}$’s IDF.
-  2. (2)
-    $\text{MS}_{\text{total}}(q_{i},d_{j})$: The total Matching Score, computed as the sum of Matching Scores from individual heads ($\text{MS}_{H_{k}}$) weighted by learned weights $\alpha_{k}$.
-Each Matching Score represents the sum of attention values from $q_{i}$ to all document tokens:
-    $\text{MS}_{\text{total}}(q_{i},d_{j})=\sum_{k=1}^{13}\alpha_{k}\cdot\text{MS}_{H_{k}}$
-  3. (3)
-    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, analogous to the product of IDF and TF in BM25.
-Since the formalization contains not only the hypothesized computation function but also $U_{0}$ and $MS$ from earlier subparts of the circuit, the linear model is an effective representation of the circuit.
-5.2. Assessing the Linear Model Fit
-Given the linear representation, we can determine whether our hypothesis holds by comparing the linear model’s performance against the cross-encoder’s actual relevance scores.
-We first train a linear regression model using our base dataset.
-As the number of coefficients in our linear regression scales quickly with each additional included query term, we apply a query length cutoff to five tokens in this initial experiment to first establish feasibility of representing the circuit in this manner.
-When performing forward passes on the samples, we extract the following for each query token: (1) the MS (Matching Score), calculated as the sum of the query token’s attention over document tokens from the 13 Matching Heads, and (2) the token’s corresponding value in $U_{0}$, obtained by applying SVD to the embedding matrix.
-These serve as the input features (“x”) for the linear regression model, while the cross-encoder’s ranking scores are used as the target values (“y”).
-We then split the data into an 80/20 train-test set to evaluate how well the linear representation of the Semantic Scoring Circuit predicts the cross-encoder’s relevance scores.
-The linear regression model achieves a high pearson correaltion (corr = 0.8157, p ¡ 0.001) with ground-truth relevance scores, showing that it effectively captures the core components of the cross-encoder’s scoring mechanism in a simplified and interpretable form.
-This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p ¡ 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components more effectively capture the signals that the cross-encoder utilizes for ranking.
-5.3. Generalization across IR datasets
-We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept.
-Now, we ask the question, how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on?
-In this section, we extend our analysis to 12 different IR datasets [^1] Due to computational resources, we use a subset of BEIR datasets  (Thakur et al., 2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains.
-and investigate various query lengths.
-Dataset.
-Since our linear model relies on fixed features based on query length, we stratify the dataset based on query length.
-To ensure sufficient samples, we aggregate samples from 12 datasets and only retain groups with more than 500 samples.
-We incorporate groups with query lengths from 1 to 22, which correponds to the 99.95th percentile of MSMARCO’s query length distribution to exclude outliers with excessively long queries that the model has rarely seen during training.
-This process results in a total of 19 groups, comprising 278194 samples, with an average group size of 14641.790 samples.
-While this approach may be unconventional for retrieval, our goal is to simply demonstrate consistency between the linear model’s relevance score and the cross-encoder to confirm the hypothesized function of the Relevance Scoring Heads.
-First, as cross-encoders are typically used for re-ranking tasks, we follow the classic re-ranking setup by first retrieving a candidate set of documents (top 10) with BM25 over all queries across the test collection of the 12 datasets.
-These candidate sets are then split into train-test groups using an 80/20 ratio.
-Experiments and Results.
-We train 22 linear regression models, one for each query length (1-22) to predict the cross-encoder’s relevance scores.
-We evaluate each model using: (1) Pearson Correlation: quantifies the correlation between predicted and actual relevance scores; (2) Spearman Rank Correlation: assesses the consistency of the ranked lists between predictions and cross-encoder outputs; (3) NDCG@10: measures ranking performance and ensures no significant performance discrepancies.
-We include BM25 and a randomized set of linear regression features as baselines.
-Table 2 shows the results, and we report median values to account for observable skewness caused by outliers in the data.
-We observe a Pearson correlation of ranking scores with a median of 0.8401 (median p ¡ 0.001), a Spearman rank correlation of 0.7619 (median p = 0.072), and an 88.4% alignment with cross-encoder performance in terms of NDCG@10.
-The experimental results confirm the hypothesized function of the Relevance Scoring Heads (§4.1).
-Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s performance shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
-Table 2. Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
 Model Pair Median Mean $\pm$ SD
@@
 Random Features 0.3562 0.3498 $\pm$ 0.2986
-6. Insights into Model Mechanisms
-A two-stage process for query-document interactions.
-Our circuit analysis reveals that query-document interactions happen in two stages, bridging previous conflicting perspectives on how BERT cross-encoders form query representations.
-In the early to mid layers (0-8), matching heads attend to document terms that are exact or “soft” matches to query terms (§4.3).
-This first stage aligns with Qiao et al.
-(2019), who find that BERT focuses on exact query term matches or close paraphrases in the document, suggesting query-dependent document representations.
-In mid to late layers (8-9), document representations are aggregated into query token representations by Contextual Query Representation heads (Section 4.2), consistent with Zhan et al.
-(2020) who argue that document tokens are largely query-independent.
-These two stages offer an explanation to previous conflicting perspectives on the role of query-document interactions: document representations initially have a query-specific relevance signal, but later, this signal is transferred to the query tokens.
-This dual-stage interaction process provides a more nuanced understanding of how query-document interactions evolve across cross-encoder layers.
-Is (semantic) BM25 the whole story?
-While the Semantic Scoring circuit bears a strong resemblance to BM25, there also may be other factors influencing relevance computation that are out of the scope of investigation for this work.
-In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022).
-However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
-Additionally, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores.
-There are a couple potential explanations for this: (1) the nonlinearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training (MacAvaney et al., 2022).
-We leave investigation of these avenues for future work.
+Table 2: Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
+Furthermore, for most heads, the TFC2 trend is preserved in the attention values.
+For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 18).
+Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes.
+The varying degrees of influence on the heads’ attention values shown in Figure 17 and 18 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects.
+Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
+Appendix E Generalization across IR datasets
+We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept.
+Now, we ask the question: how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on?
+In this section, we extend our analysis to 12 different IR datasets [^2] Due to computational resources, we use a subset of BEIR datasets Thakur et al.
+(2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains.
+and investigate various query lengths.
+Dataset.
+Since our linear model relies on fixed features based on query length, we stratify the dataset based on query length.
+To ensure sufficient samples, we aggregate samples from 12 datasets and only retain groups with more than 500 samples.
+We incorporate groups with query lengths from 1 to 22, which corresponds to the 99.95th percentile of MSMARCO’s query length distribution to exclude outliers with excessively long queries that the model has rarely seen during training.
+This process results in a total of 19 groups, comprising 278194 samples, with an average group size of 14641.790 samples.
+While this approach may be unconventional for retrieval, our goal is to simply demonstrate consistency between the linear model’s relevance score and the cross-encoder to confirm the hypothesized function of the Relevance Scoring Heads.
+First, as cross-encoders are typically used for re-ranking tasks, we follow the classic re-ranking setup by first retrieving a candidate set of documents (top 10) with BM25 over all queries across the test collection of the 12 datasets.
+These candidate sets are then split into train-test groups using an 80/20 ratio.
+Experiments and Results.
+We train 22 linear regression models, one for each query length (1-22), to predict the cross-encoder’s relevance scores.
+We evaluate each model using: (1) Pearson Correlation: quantifies the correlation between predicted and actual relevance scores; (2) Spearman Rank Correlation: assesses the consistency of the ranked lists between predictions and cross-encoder outputs; (3) NDCG@10: measures ranking effectiveness and ensures no significant effectiveness discrepancies.
+We include BM25 and a randomized set of linear regression features as baselines.
+Table 2 shows the results, and we report median values to account for observable skewness caused by outliers in the data.
+We observe a Pearson correlation of ranking scores with a median of 0.8401 (median p < 0.001), a Spearman rank correlation of 0.7619 (median p = 0.072), and an 88.4% alignment with cross-encoder effectiveness in terms of NDCG@10.
+The experimental results confirm the hypothesized function of the Relevance Scoring Heads (§4.1).
+Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s effectiveness shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
+Appendix F SVD Model Editing on Term Importance
+In §4.4, we introduce a model editing technique to dynamically control the importance signal of query tokens, and here we present encouraging evidence to use the model editing technique for group-editing the importance signals in downstream tasks.
+F.1 Mitigating Adversarial Attacks
+Figure 19: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion.
+Right: The intervention maintains a high NDCG score at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
+First, we show the potential of the model editing in mitigating unsafe documents without disrupting the cross-encoder’s ranking performance, unlike retraining or regularization methods that often degrade it.
+First, we construct a dataset using obscene and offensive words (LDNOOBW, 2015), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets.
+Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score.
+Our goal is to “erase” the effect of the dangerous token by reducing its importance.
+Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document).
+For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance.
+These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
+F.2 Parameter Efficient Fine-Tuning
+Our mechanistic analysis of the model’s relevance mechanisms reveals that term importance (IDF) is encoded in the embedding matrix.
+Since IDF determines how query terms are weighted in BM25, we hypothesize that only editing or fine-tuning the embedding matrix could provide comparable performance to fine-tuning the entire model, thus providing efficiency benefits.
+To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR Thakur et al.
+(2021).
+We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning.
+Results across three random (Figure 20 seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1).
+Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights.
+These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
+Figure 20: For random seed 15, aligned full fine-tuning achieves the best performance, and aligned embeddings only achieve performance compared to full fine-tuning.
+Appendix G Additional Discussion
 Towards a holistic approach in axiomatic analysis.
-Previous research investigating NRMs’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones (Câmara and Hauff, 2020).
+Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones Câmara and Hauff (2020).
 Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves.
@@
 Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.
-7. Conclusion
-In this work, we mechanistically uncovered the core components of relevance scoring pathway of an NRM, revealing their ability to leverage contextual representations to perform a semantic-variant of BM25.
-Our granular analysis paves way for future directions.
-For example, future work could build on our IDF editing technique (§4.4) to adjust term importance, offering more control over model behavior and leading to improvements in model robustness against adversarial attacks, personalization, and bias mitigation.
-In addition, our evidence about the centrality of IDF information in the model’s embedding matrix raises the question of how crucial non-embedding weights are for task generalization, and future work could investigate whether fine-tuning only the embedding matrix with domain-specific IDF values can achieve robust performance while being far more parameter efficient.
-Finally, in this work, we analyze the behavior of a single cross-encoder in order to deeply analyze its mechanisms, and future work should investigate their generalizability to other neural IR models.
-Overall, identifying universal components or architecture-specific mechanisms will advance interpretability and inform the design of controllable and efficient neural ranking systems for real-world applications.

### v3

@@
-Cross-Encoder Rediscovers a Semantic Variant of BM25
+Pathway to Relevance: How Cross-Encoders Implement a Semantic Variant of BM25
 Abstract
-Neural Ranking Models (NRMs) have rapidly advanced state-of-the-art performance on information retrieval tasks.
-In this work, we investigate a Cross-Encoder variant of MiniLM to determine which relevance features it computes and where they are stored.
-We find that it employs a semantic variant of the traditional BM25 in an interpretable manner, featuring localized components: (1) Transformer attention heads that compute soft term frequency while controlling for term saturation and document length effects, and (2) a low-rank component of its embedding matrix that encodes inverse document frequency information for the vocabulary.
-This suggests that the Cross-Encoder uses the same fundamental mechanisms as BM25, but further leverages their capacity to capture semantics for improved retrieval performance.
-The granular understanding lays the groundwork for model editing to enhance model transparency, addressing safety concerns, and improving scalability in training and real-world applications.
-[^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25
-Cross-Encoder Rediscovers a Semantic Variant of BM25
-Meng Lu Brown University meng_lu@brown.edu                        Catherine Chen Brown University catherine_s_chen@brown.edu
+Mechanistic interpretation has greatly contributed to a more detailed understanding of generative language models, enabling significant progress in identifying structures that implement key behaviors through interactions between internal components.
+In contrast, interpretability in information retrieval (IR) remains relatively coarse-grained, and much is still unknown as to how IR models determine whether a document is relevant to a query.
+In this work, we address this gap by mechanistically analyzing how one commonly used model, a cross-encoder, estimates relevance.
+We find that the model extracts traditional relevance signals, such as term frequency and inverse document frequency, in early-to-middle layers.
+These concepts are then combined in later layers, similar to the well-known probabilistic ranking function, BM25.
+Overall, our analysis offers a more nuanced understanding of how IR models compute relevance.
+Isolating these components lays the groundwork for future interventions that could enhance transparency, mitigate safety risks, and improve scalability.
+Pathway to Relevance:
+How Cross-Encoders Implement a Semantic Variant of BM25
+Meng Luthanks: Equal contribution. Brown University meng_lu@brown.edu          Catherine Chen11footnotemark: 1 Brown University catherine_s_chen@brown.edu
 Carsten Eickhoff University of Tübingen carsten.eickhoff@uni-tuebingen.de
@@
 Information retrieval (IR) is the subfield of NLP that aims to rank a collection of documents by their relevance to a specific query.
-Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (Robertson and Zaragoza, 2009), to estimate relevance.
+Traditional ranking strategies have long relied on probabilistic models grounded in intuitive heuristics, like BM25 (bm25), to estimate relevance.
 BM25 leverages term frequency (TF) and inverse document frequency (IDF) to rank documents effectively, achieving strong performance across various tasks.
 Its simplicity and inherent interpretability have made it a cornerstone of traditional IR systems.
-Inspired by BM25’s success, earlier neural IR models (Pang et al., 2016; Guo et al., 2016) were purposefully designed to emulate BM25’s principles.
+Inspired by BM25’s success, earlier neural IR models (pang2016text; guo2016deep) were purposefully designed to emulate BM25’s principles.
 These models incorporate explicit components for semantic TF and IDF computations, blending neural architectures with established IR heuristics to improve relevance estimation.
 However, the advent of transformer-based models revolutionized the field of IR.
-These models, trained end-to-end on large numbers of query-document pairs (Nguyen et al., 2016; Thakur et al., 2021), excel at extracting context-dependent semantic signals for ranking tasks.
-By leveraging multi-headed attention and vast parameter spaces, transformers Vaswani et al.
-(2017) capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches.
+These models, trained end-to-end on large numbers of query-document pairs (msmarco_dataset; beir), excel at extracting context-dependent semantic signals for ranking tasks.
+By leveraging multi-headed attention and vast parameter spaces, transformers vaswani_attention capture nuanced relationships between query and document terms that go beyond traditional heuristic-based approaches.
 Despite their superior performance, these models come with significant trade-offs: their complexity and lack of interpretability make it challenging to understand their internal mechanisms.
@@
 Do they rely on established IR principles such as TF and IDF, or do they draw on entirely different expressions of relevance?
-In this work, we build upon previous correlational studies (Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022; MacAvaney et al., 2022) by employing mechanistic interpretability methods (Nanda, 2022; Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2024; Wang et al., 2022; Pearl, 2022) to address these questions.
+In this work, we build upon previous correlational studies (choi_idf; zhan; formal2021white; formal2022match; macavaney2022abnirml) by employing mechanistic interpretability methods (nanda2022mechanistic; elhage2021mathematical; olsson_induction; meng_rome; wang_ioi; pearl) to address these questions.
 Concretely, we analyze a BERT-based IR model (i.e., a cross-encoder) to isolate multiple relevance signals beyond exact match TF.
-Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3).
+[^1] All code and resources are available at: https://github.com/mlu108/CrossEncoderBM25 Our main findings are: (1) We identify attention heads that extract and process BM25-like components, including a semantic version of term frequency (soft-TF), term saturation, and document length (§4.14.3).
 (2) We find evidence that IDF information largely already exists in the largest singular value of the embedding matrix and can be manipulated to later control term importance (§4.4).
 (3) We confirm that heads in later layers aggregate all these relevance signals in a BM25-style manner by defining a linear approximation of the hypothesized relevance computation and evaluating its ability to reconstruct cross-encoder’s predicted scores, confirming that our circuit captures the core mechanism of relevance computation (§5).
-Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer-based IR.
+Overall, our mechanistic analysis uncovers insights on how IR models determine relevance, paving the way for targeted model editing to boost performance, enable personalization, and mitigate bias in transformer‑based IR.
 2 Background and Related Work
@@
 (b: baseline doc, p: perturbed doc)
-TF (soft-TF) TFC1 Fang et al.
-(2004): Prefer documents with more query term occurrences.
+TF (soft-TF) TFC1 fang2004formal: Prefer documents with more query term occurrences.
 Given a baseline document, we perturb it by appending one more selected query term.
@@
 p: Quebec is a small city in Canada. Quebec.
-STMC1 Fang and Zhai (2006): Prefer documents with semantically similar terms.
+STMC1 fang2006semantic: Prefer documents with semantically similar terms.
 Given a baseline document, we append a semantically similar term - the synonym with highest embedding cosine similarity with query term from 20 candidates generated by GPT-4o.
@@
 p: Quebec is a small city in Canada. Toronto.
-IDF TDC Fang et al. (2004): Prefer documents with more discriminative query terms. N/A N/A
-k TFC2 Fang et al.
-(2004): Additional occurrences yield smaller improvements.
+IDF TDC fang2004formal: Prefer documents with more discriminative query terms. N/A N/A
+k TFC2 fang2004formal: Additional occurrences yield smaller improvements.
 We create a baseline document using GPT-4o to create five relevant sentences starting with selected query term’s pronoun and perturb by incrementally restoring the term.
@@
 p2: Quebec is in Canada. Quebec is fun.
-b LNC1 Fang et al.
-(2004): Penalize longer documents for non-relevant terms.
+b LNC1 fang2004formal: Penalize longer documents for non-relevant terms.
 Given a baseline document, we create five perturbations by sequentially appending an increasing number of random sentences from a non-relevant query in the base dataset.
@@
 Note: Since IDF is a distinct component in BM25 (unlike term saturation and document length, which are tied to TF), we use alternative methods to isolate it (§4.4).
-Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (Bruza and Huibers, 1994).
-For example, the TFC1 axiom (Fang et al., 2004) states that documents that contain more query terms should be ranked higher, and the TDC axiom (Fang et al., 2004) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores.
+Axiomatic IR constructs formal desiderata, or axioms, outlining specific properties that an effective ranking model should satisfy (axioms_bruza).
+For example, the TFC1 axiom (fang2004formal) states that documents that contain more query terms should be ranked higher, and the TDC axiom (fang2004formal) states that documents containing query terms with higher inverse document frequency (IDF) should receive higher scores.
 These axioms provide a theoretical foundation for understanding and developing ranking functions by formalizing human-interpretable and intuitive notions of relevance into mathematical constraints.
@@
 2.2 Mechanistic Interpretability for IR
-Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022; Choi et al., 2022; Zhan et al., 2020; Formal et al., 2021, 2022).
-To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (Nanda, 2022; Elhage et al., 2021; Pearl, 2022), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification Meng et al.
-(2024) and greater-than computation Hanna et al.
-(2024).
-In the context of IR, Chen et al.
-(2024) apply activation patching (Vig et al., 2020; Meng et al., 2024) to identify attention heads responsible for exact term matching.
+Despite not being explicitly trained to encode traditional relevance concepts, previous work provides correlational evidence suggesting that BERT-based IR models encode BM25-like information  (wang2021bert; rau2022different; yates2021pretrained; macavaney2022abnirml; choi_idf; zhan; formal2021white; formal2022match).
+To move beyond correlation and gain a causal understanding of how relevance is computed, we turn to mechanistic interpretability (nanda2022mechanistic; elhage2021mathematical; pearl), which has been instrumental in uncovering how transformer-based NLP models perform certain tasks, such as indirect object identification meng_rome and greater-than computation hanna2024does.
+In the context of IR, chen_axiomatic apply activation patching (Vig; meng_rome) to identify attention heads responsible for exact term matching.
 We build on this work by decomposing BM25 into intuitive IR axioms and constructing diagnostic datasets to localize each component.
-Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023).
+Rather than focusing solely on exact matches, we extend the analysis to semantic matches and trace how these components interact to compute the final relevance score using path patching (wang_ioi; goldowskydill_ppatching).
 This work is the first to causally uncover an internal circuit for the relevance estimation mechanism of neural IR architectures.
@@
 Given input x in the format: <CLS> query <SEP> document <SEP>, the model is trained to optimize a binary classification task where the <CLS> token is passed to a classifier to determine whether the provided query is relevant to the document.
-In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (Face, 2025), a BERT-based cross-encoder, for its high performance on common IR benchmarks (Nguyen et al., 2016; Craswell et al., 2020).
+In this work, we choose to examine ms-marco-MiniLM-L-12-v2 (minilm_huggingface), a BERT-based cross-encoder, for its high performance on common IR benchmarks (msmarco_dataset; trec_dataset).
 3.2 Diagnostic Datasets Construction
 To investigate if and how the model implements BM25, we map BM25 components to IR axioms (§2.1) and construct diagnostic datasets that isolate individual axiomatic components.
-We build off of Chen et al.
-(2024), who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (Nguyen et al., 2016), consisting of 10k web search query-document pairs.
+We build off of chen_axiomatic, who create a diagnostic dataset for analyzing TFC1 using MS-MARCO (msmarco_dataset), consisting of 10k web search query-document pairs.
 To create a diagnostic dataset for the remaining axioms, we perturb each query-document pair in this base dataset following their formal axiomatic definition (Table 1).
@@
 Our perturbation strategies and examples of input pairs are shown in Table 1.
-Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (Rau and Kamps, 2022).
+Additionally, we expand the analysis on TF to soft-TF to include semantic matches, motivated by findings that neural models often score documents highly even in the absence of exact lexical matches (rau2022different).
 3.3 Path Patching on Diagnostic Datasets
@@
 The path patching algorithm proceeds iteratively in a backward fashion: it first identifies components directly responsible for changes in the output logits, then recursively traces upstream to uncover the full set of causally relevant components.
-This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix §A).
+This backward search reveals the components through which specific information, introduced by the perturbation, flows (more details in Appendix C).
 4 Semantic Scoring Circuit
 Figure 1: Overview of relevance mechanisms in the model.
-In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (Figure 1):
+The model first jointly analyzes query and document tokens to identify matching terms (exact and semantic), then contextualizes each query term within the full query, and finally calculates a relevance score by weighting query terms by their importance (IDF) and aggregating them, similar to BM25.
+In the course of this section, we will identify and localize the following components in the model’s Semantic Scoring Circuit (overview in Figure 1, detailed walk-through example in Figure 10):
   • Matching Heads locate exact and semantic term matches in the document for each query term, while also encoding term saturation and document length signals.
@@
 To identify the components that directly transmit soft-TF signals to the relevance scores, we path patch to the logits using the TFC1 and STMC1 diagnostic datasets.
-We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that the same heads produce similar effects on delivering soft-TF information into the final logits (see Appendix 9 for detailed patching results).
-Figure 2 displays the heads with the highest patching effects, which are consistent across both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10.
+We observe a high Pearson correlation between the patching effects of TFC1 and STMC1 (r = 0.99, p < 0.001), indicating that path patching on both datasets identifies essentially the same set of influential heads.
+This suggests that the model treats exact and semantic matches similarly in the final relevance computation (see Figure 15 in Appendix D for detailed patching results).
+Figure 2 displays the heads with the highest patching effects that exceed the top 30% of causal importance for both TFC1 and STMC1: 10.1 (Layer 10, Head 1), 10.4, 10.7, and 10.10.
 Notably, these heads contribute a larger change in relevance scores for TFC1 compared to STMC1, suggesting that the model may prioritize exact matches over semantic ones.
@@
 Qualitative analysis reveals that the [CLS] token selectively attends to query tokens, indicating the movement of soft-TF signals into its representation to prepare for the final relevance score in the model’s classification head.
-We find the heads distribute attention to different parts of the query.
+We find that the heads distribute attention to different parts of the query.
 Specifically, 10.1 focuses on high-IDF query terms, 10.10 on low-IDF, and 10.7 and 10.4 on mid-to-high-IDF (Figure 3).
@@
 We path patch to the Relevance Scoring Heads’ value vectors to identify which upstream components transmit soft-TF signals.
-On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix C).
+On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix D).
 4.2 Query Contextualization Heads
@@
 Further study of this learned contextualization is left for future work.
-Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (Appendix C).
+Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (more details in Appendix D).
 4.3 Matching Heads
@@
 As shown in Figure 4, head 1.7 exhibits this behavior clearly: both query and document tokens strongly attend to duplicated terms and, to a lesser extent, to semantically related terms.
-We compute the Pearson correlation between attention scores and semantic similarity for each query-document token pair, with semantic similarity measured as the cosine similarity between token embedding vectors.
-Matching Heads on average exhibit a significantly stronger correlation (r = 0.500, p < 0.001) than other heads (r = 0.132).
+To quantitatively verify this behavior, we compute the Pearson correlation between attention weights and semantic similarity for each query–document token pair.
+If a head indeed matches semantically similar words, its attention from a query token $q_{i}$ to a document token $d_{j}$ should increase with $\cos(\mathbf{q}_{i},\mathbf{d}_{j})$.
+On average, Matching Heads exhibit a substantially stronger correlation ($r=0.500$, $p<0.001$) compared to other heads ($r=0.132$).
 This evidence suggests that Matching Heads employ an attention mechanism that scales proportionally with semantic proximity: higher semantic similarity corresponds to higher attention scores.
@@
 Because each dataset varies only in one factor, we track each Matching Head’s average attention across these controlled groups.
-Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (see Appendix D): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length.
+Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (additional details in Appendix E): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length.
 These findings suggest that Matching Heads integrate soft-TF with saturation and document-length signals rather than operating in isolation.
@@
 Important Upstream Components.
-Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix C.2), confirming they generate soft-TF based on semantic similarity from the embeddings.
+Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix D.2), confirming they generate soft-TF based on semantic similarity from the embeddings.
 This completes the backward tracing for the soft-TF circuit.
@@
 As shown in §4.1, IDF similarly emerges as a key signal for the Relevance Scoring Heads in our model.
-While Choi et al.
-(2022) provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
+While choi_idf provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
 SVD allows us to decompose a matrix into a sum of orthogonal rank-1 components, outer products of a column and row vector, ordered by their contribution to the overall matrix.
@@
 By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the embedding matrix.
-We find that the top singular vector $U_{0}$ is highly correlated (r = $-71.36\%$) with MS-MARCO IDF values Nguyen et al.
-(2016), the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
+We find that the top singular vector $U_{0}$ is highly correlated (r = $71.36\%$) with MS-MARCO IDF values msmarco_dataset, the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
 Figure 5: $U_{0}$ editing example (transposed for visualization).
@@
 To test this, we design an experiment with controlled, minimal examples.
-Given a query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”).
+Given each query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”).
 Then, we can edit the IDF of tok1 and measure the effect on the relevance scores of doc1 and doc2.
@@
 Increasing tok1’s IDF raises doc1’s score more than doc2’s.
-Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
-Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values (§4.1).
+Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight non-monotonic deviations suggesting optimal editing windows.
+Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values as hypothesized in §4.1 (see additional details about IDF and Relevance Scoring Heads’ attention pattern correlation in Appendix F.1).
 Comparison to BM25.
-This completes the BM25 component set: IDF embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF.
+This completes the BM25 component set: the embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF.
 This is similar to BM25, where each query term’s TF is weighted by its IDF to prioritize informative terms.
@@
 To validate whether Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function.
-Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
+Next, we evaluate this linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
 5.1 Formalizing the Hypothesized Function
-In §4.1, we hypothesize that Relevance Scoring Heads appear to compute a summation of Matching Scores weighted by IDF.
+In §4.1, we hypothesize that Relevance Scoring Heads compute a summation of Matching Scores weighted by IDF.
 The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function.
-If this hypothesis holds, that these components interact in a BM25-like manner, then the semantic scoring circuit can be expressed as a linear function:
-[…]
+If this hypothesis that these components interact in a BM25-like manner holds, then the Semantic Scoring Circuit can be expressed as a linear function:
+$\sum_{i=1}^{N}$ $\;\text{linear\_combo}\Bigl(-U_{0}(q_{i}),\;\text{MS}_{\mathrm{total}}(q_{i},d_{j}),$ (2)
 $\quad-\,U_{0}(q_{i})\,\cdot\,\text{MS}_{\mathrm{total}}(q_{i},d_{j})\Bigr)$
@@
   3. 3.
-    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
+    $U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
 By incorporating both the hypothesized computation function and the earlier components $U_{0}$ and $MS$, the linear model effectively represents the hypothesized circuit.
@@
 The linear regression model achieves a high Pearson correlation (r = 0.8157, p < 0.001) with ground-truth relevance scores, showing that it captures the core of the cross-encoder’s scoring mechanism in a simplified and interpretable form.
-This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking.
-The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix §E) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
+This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates that the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking.
+The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix G) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
 6 Discussion
-A two-stage process for relevance computation.
-Similar to Tenney et al.
-(2019), who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process.
-In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (Qiao et al., 2019).
-In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with Zhan et al.
-(2020), who argue that document tokens are largely query-independent.
+6.1 A Two-Stage Process for Relevance Computation
+Similar to tenney2019bert, who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process.
+In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (qiao2019understanding).
+In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with zhan, who argue that document tokens are largely query-independent.
 This dual-stage mechanism reconciles these two conflicting viewpoints by showing how query-specific signals migrate from document to query representations across layers and provides a more nuanced understanding of how relevance is computed.
-Controllable IR.
+6.2 Potential Downstream Applications
 In §4.4, we show how model editing can be used to upscale or downscale term importance by modifying the encoded IDF values in the embeddings, effectively introducing “tuning dials” for fine-grained control over model behavior.
-This opens the door to personalization, where models could be adapted to user- or domain-specific preferences without full retraining.
-Beyond personalization, these insights enable two additional downstream applications (preliminary experiments in Appendix §F).
+These insights enable two potential downstream applications.
+6.3 Mitigating Adversarial Attacks
+Figure 7: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion.
+Right: The intervention maintains a high NDCG at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
 First, targeted model editing could mitigate unsafe or adversarial content without impairing the model’s ranking effectiveness.
+As an initial test, we construct a dataset using obscene and offensive words (ldnoobw_badwords), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets.
+Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score.
+Our goal is to “erase” the effect of the dangerous token by reducing its importance.
+Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document).
+For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance.
+These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
+6.4 Parameter Efficient Fine-Tuning
+Figure 8: For random seed 15, aligned full fine-tuning achieves the best performance, and aligned embeddings only achieve performance compared to full fine-tuning.
 Second, aligning internal representations with ground-truth IDF scores could serve as a powerful initialization strategy for fine-tuning, leading to more efficient retrieval pipelines.
+In other words, rather than updating the entire parameter space, fine-tuning the embedding matrix to better encode IDF values aligned with the retrieval domain may suffice.
+To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR beir.
+We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning.
+The results in Figure 8 across three random seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1).
+Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights.
+These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
 7 Conclusion
@@
 Additionally, while the Semantic Scoring circuit bears a strong resemblance to BM25, there may also be other factors influencing relevance computation that are out of the scope of investigation for this work.
-In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022).
+In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that neural IR encode a BM25-like relevance signal (wang2021bert; rau2022different; yates2021pretrained; macavaney2022abnirml).
 However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
 Furthermore, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores.
-There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training MacAvaney et al.
-(2022).
+There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training macavaney2022abnirml.
 We leave the investigation of these avenues for future work.
 Finally, although we intend for our analysis to inform future work on efficiency, transparency, and performance improvements, like most research, it is possible that a malicious actor could apply these insights to produce IR models that are robust to adversarial use cases.
-Acknowledgment
+Acknowledgments
 We thank the Health NLP Lab for valuable suggestions and discussions that improved this work.
-Appendix A Detailed Path Patching Methodology
+Appendix A Activation Patching on Attention Heads
+Prior to path patching, we conduct an exploratory analysis to compare the similarity between exact and semantic matches.
+Figure 9 shows that the patching effect has a high correlation of 0.96, suggesting that the model employs highly similar components for both exact and soft matches.
+These initial results on the indirect effect of these components provide a foundation to investigate their direct effects further with path patching.
+Figure 9: Activation Patching Results for TFC and STMC diagnostic datasets show a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
+Appendix B Example of Semantic Scoring Circuit
+Figure 10: Example walkthrough of the hypothesized circuit.
+We find that the cross-encoder rediscovers a semantic variant of BM25.
+Specifically, Matching Heads, Contextual Query Representation Heads, and the embedding matrix compute, process and send BM25-like components’ information to the Relevance Scoring Heads which finally compute the relevance score in a BM25-like manner.
+The trapezoid represents the residual stream representation for each token position.
+Figure 10 shows an example of the hypothesized semantic scoring circuit.
+In the figure, the model receives an example input: “[CLS] Where is Wellesley [SEP] Wellesley is a college in Massachusetts. [SEP]“.
+(1) Matching Heads: These heads generate $MS(\texttt{where})$, $MS(\texttt{is})$, and $MS(\texttt{wellesley})$ that mainly capture query term soft-TF values.
+For example, the query token wellesley strongly attends to wellesley in the document, and more weakly to related terms like wellesley and college.
+This attention pattern forms a Matching Score $MS(\texttt{wellesley})$ that also reflects document length normalization (the longer the document, the smaller the soft $MS(\texttt{wellesley})$) and term saturation effects (the more wellesley exists in the document, the less additional gain in $MS$).
+(2) Query Contextualization Heads: These heads redistribute the soft-TF signal from key content words to surrounding query tokens.
+For instance, where and is may receive part of wellesley’s soft-TF signal, allowing the model to amplify high-IDF query terms’ soft-TF signals.
+(3) IDF storage: The IDF of all query terms is stored in a dominant low-rank vector of the model’s embedding matrix.
+(4) Relevance Scoring Heads: Finally, the model aggregates the reweighted soft-TF signals across query tokens using a mechanism similar to BM25.
+Appendix C Detailed Path Patching Methodology
 Figure 11: Path patching methodology.
@@
 This algorithm is then iteratively repeated for each important upstream component.
-For more information, we refer the reader to Wang et al.
-(2022) and Goldowsky-Dill et al.
-(2023).
+For more information, we refer the reader to wang_ioi and goldowskydill_ppatching.
 We begin by applying path patching to track the path of term-matching, the most fundamental component of BM25.
 To carry this out, we use the TFC1 and STMC1 diagnostic datasets to identify the components responsible for encoding the relevance signal of an additional query term in a document, providing the foundation for our analysis of model behavior.
-Appendix B Activation Patching on Attention Heads
-Prior to path patching, we conduct an exploratory analysis to compare the similarity between exact and semantic matches.
-Figure 8 shows that the patching effect has a high correlation of 0.96, suggesting that the model employs highly similar components for both exact and soft matches.
-These initial results on the indirect effect of these components provide a foundation to investigate their direct effects further with path patching.
-Appendix C Detailed Path Patching Results
-C.1 Path Patching to Final Logits and Intermediate Heads
+Appendix D Detailed Path Patching Results
+D.1 Path Patching to Final Logits and Intermediate Heads
 We perform path patching on 2000 random query and document pairs from the STMC1 and TFC1 diagnostic datasets and collect the results.
@@
 Finally, by patching to the Query Contextualization Heads, we identify the shared Matching Heads (Figure 18).
-Figure 8: Activation Patching Results for TFC and STMC diagnostic datasets shows a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
-Figure 9: Path Patching results from all attention heads to the final logits.
-From this, we identify that 10.1, 10.4, 10.7, 10.19, those causing > 30 % increase in ranking score, as most significant heads and focus our analysis on these four heads.
-Figure 10: Path Patching results from upstream attention heads to Relevance Scoring Heads.
-Figure 11: Path Patching results from upstream attention heads to Relevance Scoring Heads.
-Figure 12: Path Patching results from upstream attention heads to Query Contextualization heads.
-C.2 Path Patching to Matching Heads
-We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 13, 14, 15, 16).
+All resulting heatmaps are located at the end of the appendix for clearer flow for the other appendix sections.
+D.2 Path Patching to Matching Heads
+We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 19, 20, 21, 22).
 Path patching to the Matching Heads, aimed at tracking the flow of soft-TF information, reveals that at the beginning of the residual stream at layer 0, which corresponds to the embeddings, has the most direct influence on the Matching Heads.
@@
 Thus, we conclude our tracking of soft-TF information, with the Matching Heads identified as the most upstream heads.
-Figure 13: TFC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
-Figure 14: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
-Figure 15: STMC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
-Figure 16: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
-Appendix D Additional Details on Matching Heads and Additional Signals
-In this section, we provide additional details on the analysis presented in §4.3.
-Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence.
-The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, proving that these signals are discrete rather than binary.
-To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset.
-Figure 17 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals.
-This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
 Figure 12: For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases.
@@
 Figure 13: Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
+Appendix E Additional Details on Matching Heads and Additional Signals
+In this section, we provide additional details on the analysis presented in §4.3.
 Matching Score Contains Term Saturation and Document Length Signals.
@@
 To examine whether these attention values capture term saturation or document length effects, we analyze their behavior under using two diagnostic datasets: (1) TFC2 (increasing occurrences of a query term), and (2) LNC1 dataset (increasing number of irrelevant sentences to simulate the isolated effect of longer document length).
+Saturation Test (TFC2).
 We define $x_{\text{selected}}$ as the selected query term, which is the duplicate term in TFC2 samples shown in Table 1.
@@
 For each case, we calculate the normalized sum of attention values from all query tokens to $x_{\text{selected}}$ and $x_{\text{others}}$, representing the “total semantic match” of the matched token or the average across unmatched tokens, respectively.
+We hypothesize that if the Matching Score contains the term saturation signal, then given more duplicate query terms (TFC2), the summed attention for that term will rise sharply at first and then plateau.
 As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 12).
@@
 Table 2: Comparison of Pearson, Spearman Rank, and NDCG@10 for SemanticBM and Baselines.
-Furthermore, for most heads, the TFC2 trend is preserved in the attention values.
-For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 18).
+Length Test (LNC1).
+We hypothesize that if the Matching Score contains the term saturation signal, given irrelevant sentences appended (LNC1), the overall attention to the query’s matches will fall.
+As expected, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 13).
 Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes.
-The varying degrees of influence on the heads’ attention values shown in Figure 17 and 18 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects.
+The varying degrees of influence on the heads’ attention values shown in Figure 12 and 13 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects.
 Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
-Appendix E Generalization across IR datasets
+Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence.
+The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, suggesting that these signals are discrete rather than binary.
+To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset.
+Figure 12 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals.
+This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
+Appendix F Additional Details on BM25-like Computation
+In this section, we provide additional details on the Relevance Scoring Heads and the Semantic Scoring Hypothesis in §4.1 and §5.
+F.1 Correlation Between IDF Values and Attention Distribution
+Figure 14: Top: Changes in attention at 10.1 from [CLS] to tok1 when increasing tok1’s IDF at different scales, averaged across all samples.
+Increasing tok1’s IDF raises attention from [CLS] to tok1, thereby increasing the final score.
+Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
+In Section 4.1, we hypothesize that Relevance Scoring Heads combine IDF and soft-TF information to compute final relevance scores in a BM25-like manner.
+Specifically: (A) increasing or decreasing a token’s IDF leads to (B) the Relevance Scoring Head 10.1222We focus on the attention pattern of 10.1 because it is most positively correlated with IDF (§4.1) and has highest path-patching value.
+allocating more or less attention from [CLS] to the token, which (C) increases or decreases to the weighting of the token’s soft-TF and thereby the final relevance score.
+A $\rightarrow$ B.
+[…]
+A $\rightarrow$ B $\rightarrow$ C.
+In Section 4.4, we modulate IDF by editing the first SVD component of the embedding matrix.
+Figure 6 shows that scaling a token’s IDF upward or downward produces a monotonic change in the relevance score.
+Taken together with Section 5 (which shows that Relevance Scoring Heads perform BM25-style computations on natural IR datasets), these results provide stronger causal evidence for the Semantic Scoring hypothesis.
+F.2 Independence between IDF and Soft-TF
+In this subsection, we check whether term-frequency (TF) and inverse document frequency (IDF) factors are fully separable.
+Low correlation between IDF Soft-TF.
+If soft-TF scores are already IDF-weighted, we would expect their one-dimensional values to correlate at least mildly with $U0$.
+However, across all 13 Matching Heads, the average Pearson correlation between $U0$ and each head’s Matching-Score is low ($r=0.143$, $p<0.01$), suggesting minimal dependence.
+Relevance Scoring Heads Do Not Inherit IDF Information From Matching Heads.
+Path patching from the Matching Heads to the query vectors of the Relevance Scoring Heads produces almost no change in output score.
+This indicates that the soft-TF signals produced by the Matching Heads do not strongly convey IDF information, leaving the Relevance Scoring Heads without sufficient signal to prioritize which tokens should receive higher weight.
+Appendix G Generalization Across IR Datasets
 We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept.
 Now, we ask the question: how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on?
-In this section, we extend our analysis to 12 different IR datasets [^2] Due to computational resources, we use a subset of BEIR datasets Thakur et al.
-(2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains.
+In this section, we extend our analysis to 12 different IR datasets [^3] Due to computational resources, we use a subset of BEIR datasets beir (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains.
 and investigate various query lengths.
@@
 Since this linear model summarizes the whole circuit as it is structured to incorporate both the computation and the necessary components, the high correlation with the cross-encoder’s effectiveness shows that our circuit understanding has captured the core part of the cross-encoder’s relevance ranking mechanism.
-Appendix F SVD Model Editing on Term Importance
-In §4.4, we introduce a model editing technique to dynamically control the importance signal of query tokens, and here we present encouraging evidence to use the model editing technique for group-editing the importance signals in downstream tasks.
-F.1 Mitigating Adversarial Attacks
-Figure 19: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion.
-Right: The intervention maintains a high NDCG score at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
-First, we show the potential of the model editing in mitigating unsafe documents without disrupting the cross-encoder’s ranking performance, unlike retraining or regularization methods that often degrade it.
-First, we construct a dataset using obscene and offensive words (LDNOOBW, 2015), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets.
-Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score.
-Our goal is to “erase” the effect of the dangerous token by reducing its importance.
-Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document).
-For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance.
-These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
-F.2 Parameter Efficient Fine-Tuning
-Our mechanistic analysis of the model’s relevance mechanisms reveals that term importance (IDF) is encoded in the embedding matrix.
-Since IDF determines how query terms are weighted in BM25, we hypothesize that only editing or fine-tuning the embedding matrix could provide comparable performance to fine-tuning the entire model, thus providing efficiency benefits.
-To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR Thakur et al.
-(2021).
-We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning.
-Results across three random (Figure 20 seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1).
-Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights.
-These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
-Figure 20: For random seed 15, aligned full fine-tuning achieves the best performance, and aligned embeddings only achieve performance compared to full fine-tuning.
-Appendix G Additional Discussion
+Appendix H Additional Discussion
 Towards a holistic approach in axiomatic analysis.
-Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones Câmara and Hauff (2020).
+Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones camara2020diagnosing.
 Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves.
@@
 Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.
+Figure 15: Path Patching results from all attention heads to the final logits.
+From this, we identify that 10.1, 10.4, 10.7, 10.19, those causing > 30 % increase in ranking score, as most significant heads and focus our analysis on these four heads.
+Figure 16: Path Patching results from upstream attention heads to Relevance Scoring Heads.
+Figure 17: Path Patching results from upstream attention heads to Relevance Scoring Heads.
+Figure 18: Path Patching results from upstream attention heads to Query Contextualization heads.
+Figure 19: TFC1 path patching into Matching Heads (layers 8.8, 8.1, 8.0, 7.9, 6.5, 5.9).
+The residual stream values are patched, with lighter regions indicating stronger path-patching contributions (i.e., where the TFC1 signal flows), and darker regions indicating weaker contributions.
+Figure 20: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
+Figure 21: STMC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
+Figure 22: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)